### v1

# Chern Character and Fermi point

Abstract.
This paper expresses the Chern character for topological $K$-theory based on the formulation of the family of Fredholm operators, by using the points at which the Fredholm operator becomes singular (Fermi points). In particular, we explain that the odd Chern character can be thought of as a generalization of the spectral flow. As applications, we give elementary proofs of the evenness of the edge index and the bulk-edge correspondence for four-dimensional topological insulators with time-reversal symmetry of class AI.
1. Introduction
1.1. Topological $K$-theory and spectral flow
Topological $K$-theory, which was introduced by Atiyah and Hirzebruch, is constructed by isomorphism classes of complex vector bundles [1]. The construction by using complex vectors is a formulation within the scope of finite-dimensional linear algebra, but there is also a construction via infinite dimensions. That is, the formulation of topological $K$-theory by using families of Fredholm operators acting on infinite-dimensional Hilbert spaces with a symmetry of Clifford algebras [3]. The advantage of the Fredholm operator formulation is that it is suitable for twisted $K$-theory, which is a variant of topological $K$-theory. Moreover, $K$-theory by using Fredholm operators appears naturally in a description of the edge states of topological insulators.
Since the Fredholm operator formulation involves infinite dimensions, it is generally difficult to directly calculate the $K$-group. However, there exists a feasible invariant called the spectral flow. The spectral flow becomes a complete invariant of the odd $K$-group of the circle $S^{1}$ [5]. That is
$\text{sf}\colon K^{1}(S^{1})\cong\mathbb{Z}.$
[…]
$F=\{x\in X\mid 0\in\text{Spec}(A_{x})\},$
where $\text{Spec}(A_{x})$ denotes the spectrum of $A_{x}$. We refer to a point $x\in F$ at which a sign coordinate can be defined as a Fermi point. With each Fermi point $x\in F$, we associate a sign, denoted as $\operatorname{sign}(x)\in\{\pm 1\}$, which corresponds to the contribution of $x$ to the generalized spectral flow.
We note that the notion of a Fermi point used in this paper differs from its conventional usage in condensed matter physics. Typically, the Fermi set refers to the set of momenta with a prescribed energy level, and isolated points in this set are referred to as Fermi points. In contrast, in this paper, we use the term Fermi point to refer specifically to those isolated points in the set $F$ at which a sign coordinate can be defined. Thus, the term is used in a way that deviates from its standard meaning in physics.
1.2. Main result
We present the main theorems of this paper, which express the Chern character in terms of Fermi points. Here, $\mathcal{F}_{0}(\hat{\mathcal{H}})$ denotes the space of Fredholm operators of degree 1 acting on a $\mathbb{Z}_{2}$-graded Hilbert space $\hat{\mathcal{H}}$, while $Fred_{sa}$ denotes the space of self-adjoint Fredholm operators on an ungraded Hilbert space $\mathcal{H}$, and $Fred_{sa}^{1}$ denotes the non-contractible path-connected component of $Fred_{sa}$. The Chern characters $\operatorname{ch}_{k}$ and $\operatorname{ch}_{k+\frac{1}{2}}$ are invariants associated with the $K$-theory classes in even and odd degrees, respectively. Precise definitions of these terms will be provided in later sections.
Theorem 1.1 (The Main Theorem for Even Degrees).
[…]
  1. (1)
    When each point $x\in F$ is a Fermi point, $\sum_{x\in F}\operatorname{sign}(x)$ depends only on $[\hat{A}]\in K^{0}(X)$.
  2. (2)
    Under the assumption of $(1)$, the following equality holds.
    $\int_{X}\operatorname{ch}_{k}([\hat{A}])=\sum_{x\in F}\operatorname{sign}(x)\in\mathbb{Z}.$
Theorem 1.2 (The Main Theorem for Odd Degrees).
Let $k$ be a non-negative integer. Let $X$ be a compact, oriented, $12k+$ 1-dimensional differentiable manifold, and  $A\colon X\to Fred_{sa}^{1}(\mathcal{H})$ a continuous map. Then, the following holds.
  1. (1)
    When each point $x\in F$ is a Fermi point, $\sum_{x\in F}\operatorname{sign}(x)$ depends only on $[\hat{A}]\in K^{1}(X)$.
  2. (2)
    Under the assumption of $(1)$, the following equality holds.
    $\int_{X}\operatorname{ch}_{k+\frac{1}{2}}([A])=\sum_{x\in F}\operatorname{sign}(x)\in\mathbb{Z}.$
In particular, when considering the case of odd degrees with $X=S^{1}$, the odd Chern character $\operatorname{ch}_{\frac{1}{2}}([A])$ integrated over $S^{1}$ coincides with the spectral flow up to a sign difference.
1.3. Applications to Topological Insulators
Topological insulators are quantum systems characterized by insulating bulk states and conducting edge states. A key feature of these systems is the bulk-edge correspondence, which relates topological invariants of the bulk to those of the edge, a phenomenon that has been formulated in various settings (e.g., [7], [8], [9], [10], [16], [18], [21]).
In this paper, we focus on four-dimensional topological insulators with time-reversal symmetry of class AI, corresponding to spinless or integer-spin quantum systems invariant under time reversal (class AI). These systems are described by a bulk Hamiltonian $\hat{H}\colon T^{4}\to\operatorname{Herm}(\mathbb{C}^{r})^{*}$, where $T^{4}$ is the four-dimensional torus representing the wave number space, and $\operatorname{Herm}(\mathbb{C}^{r})^{*}$ denotes the set of invertible $r\times r$ Hermitian matrices. The edge Hamiltonian $\hat{H}^{\#}={\{\hat{H}^{\#}(k)\}}_{k\in T^{3}}$, parameterized by the three-dimensional torus $T^{3}$, is constructed from $\hat{H}$ as a family of Toeplitz operators [12].
For four-dimensional class AI topological insulators, we define the bulk index as the second Chern number $c_{2}(\hat{H})$ and the edge index as the odd Chern character $\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\#}])$. Under class AI symmetry, the edge index is known to be an even integer, and the bulk-edge correspondence is formulated as $c_{2}(\hat{H})=\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\#}])$.
By applying our main theorems, which express the Chern character in terms of Fermi points, we obtained the following results for four-dimensional class AI topological insulators:
  1. (1)
    Assuming class AI time-reversal symmetry and a stronger condition, the edge index satisfies $\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\#}])\in 2\mathbb{Z}$. (Proposition 4.22)
  2. (2)
    The bulk-edge correspondence holds, i.e., $c_{2}(\hat{H})=\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\#}])$, for insulators defined on the torus $T^{4}$ with class AI symmetry. (Theorem 5.7.)
1.4. Outline
This paper is organized as follows: In §2, we summarize the Clifford algebras and their representations necessary for the formulation of $K$-groups via Fredholm operators. In §3, we discuss the $K$-theory associated with families of Fredholm operators. In §§3.1, we first define the separable infinite-dimensional Hilbert space $\mathcal{H}_{n}$ endowed with a symmetry of the Clifford algebra $Cl_{n}$. Subsequently, we provide the definition of the $K$-group based on families of Fredholm operators acting on $\mathcal{H}_{n}$. Next, in §§3.2, we describe the suspension isomorphism of $K$-groups in the context of the Fredholm operator formulations, as developed by Atiyah and Singer [3]. As an application of the suspension isomorphism, we explicitly characterize the generators of $K^{n}(D^{n},\partial D^{n})$. Since detailed descriptions of the generators of $K^{n}(D^{n},\partial D^{n})$ in the Fredholm operator formulation are rarely found in the literature, we elaborate on this point. In §§3.3, we define the Chern character by using Fredholm operators and introduce the push-forward map along a point insertion. In §§4.1, we explain the finite-dimensional approximation of Fredholm operators. By using this approximation, we define a sign coordinate and a Fermi point in §§4.2. In §§4.3, we provide the proof of the main theorem. In §§4.4, we present specific computational examples derived from the main theorem. In §§4.5, we state the assumptions of Application $(1)$ ((((Proposition 4.22)))) and provide its proof. In §§5.1, we introduce the tools necessary for proving the bulk-edge correspondence. Specifically, we define real vector bundles in the sense of Atiyah and $\mathbb{Z}_{2}$-equivariant homotopy sets. Building on this, in §§5.2, we prove Application $(2)$, namely the bulk-edge correspondence for four-dimensional topological insulators in class AI. Finally, in §6, we detail the spectral analysis of local models used in the specific computational examples of §§4.4 and the proof of the bulk-edge correspondence in §§5.2.
Throughout this paper, we denote the identity map on the set $X$ by $1_{X}$. Additionally, we denote the interior of the topological space $X$ by $\text{int}(X)$.
Acknowledgements.
The author would like to thank my supervisor, Kiyonori Gomi. The author is grateful to Yuki Matsuoka for helpful discussions and comments. This work was supported by JST SPRING, Japan Grant Number JPMJSP2180.
2. Clifford algebra and its representation
We summarize the basics about Clifford algebra. The references include [2], [24], and [13].
Definition 2.1.
Let $n\in\mathbb{N}$ be a natural number and $\delta_{ij}$ denote the Kronecker delta. The Clifford algebra $Cl_{n}$ is an associative algebra over $\mathbb{R}$ with generators $e_{1},...,e_{n}$ which satisfy the following relations:
$e_{i}e_{j}+e_{j}e_{i}=2\delta_{ij}.$
Let $Cl_{0}=\mathbb{R}.$
Definition 2.2.
  1. (1)
    Let $\hat{V}$ be a vector space, and let $V^{0},V^{1}\subset\hat{V}$ be subspaces. In this case, if $\hat{V}$ admits a direct sum decomposition $\hat{V}=V^{0}\oplus V^{1}$ by $V^{0}$ and $V^{1}$, then $\hat{V}$ is said to be $\mathbb{Z}_{2}$-graded by $V^{0}$ and $V^{1}$. A vector space $\hat{V}$ that is $\mathbb{Z}_{2}$-graded is called a $\mathbb{Z}_{2}$-graded vector space.
  2. (2)
    A degree-0 linear map $\hat{f}:\hat{A}\to\hat{B}$ between two $\mathbb{Z}_{2}$-graded vector spaces $\hat{A}$ and $\hat{B}$ is defined as a linear map that satisfies $\hat{f}(A^{0})\subset B^{0}$ and $\hat{f}(A^{1})\subset B^{1}$.
  3. (3)
    A degree-1 linear map $\hat{f}:\hat{A}\to\hat{B}$ between two $\mathbb{Z}_{2}$-graded vector spaces $\hat{A}$ and $\hat{B}$ is defined as a linear map that satisfies $\hat{f}(A^{0})\subset B^{1}$ and $\hat{f}(A^{1})\subset B^{0}$.
In general, a $\mathbb{Z}_{2}$-graded vector space $\hat{V}$ is the direct sum of the parts of degree 00 and $1$: $\hat{V}=V^{0}\oplus V^{1}$. This is equivalent to considering the eigendecomposition of the involution $\hat{\epsilon}\colon\hat{V}\to\hat{V}$ by the eigenspaces $V^{i}=\ker(\hat{\epsilon}-(-1)^{i})$. In this case, the linear mapping $\hat{f}\colon\hat{V}\to\hat{V}$ is of degree 00 (resp. $1$), which is equivalent to the commutativity (or anticommutativity) of $\hat{f}$ and $\hat{\epsilon}$.
Definition 2.3.
Let $n\in\mathbb{N}$ be a natural number.
  1. (1)
    A pair $(V,\gamma)$ is an ungraded (complex) representation of $Cl_{n}$ if $V$ is a complex vector space and $\gamma=(\gamma_{1},...,\gamma_{n})$ are complex linear maps on $V$ such that
    $\gamma_{i}\circ\gamma_{j}+\gamma_{j}\circ\gamma_{i}=2\delta_{ij}1_{V},$
    where $1_{V}$ is the identity map on $V$.
  2. (2)
    A pair $(\hat{V},\hat{\gamma})$ is a $\mathbb{Z}_{2}$-graded (complex) representation of $Cl_{n}$ if $\hat{V}$ is $\mathbb{Z}_{2}$-graded complex vector space and $\hat{\gamma}=(\hat{\gamma_{1}},...,\hat{\gamma_{n}})$ are complex linear maps of degree $1$ on $\hat{V}$ such that
    $\hat{\gamma_{i}}\circ\hat{\gamma_{j}}+\hat{\gamma_{j}}\circ\hat{\gamma_{i}}=2\delta_{ij}1_{\hat{V}},$
    where $1_{\hat{V}}$ is the identity map on $\hat{V}$.
Here, $V$ and $\hat{V}$ are called representation spaces. In either case, a representation that has no nontrivial subrepresentation is called an irreducible representation.
Remark 2.4.
Without loss of generality, we can assume that $V$ and $\hat{V}$ have Hermitian inner products and that $\gamma_{i}$ and $\hat{\gamma_{i}}$ are unitary operators. In that case, $\gamma_{i}$ and $\hat{\gamma_{i}}$ are self-adjoint, since $\gamma_{i}^{2}=1$ and $\hat{\gamma_{i}}^{2}=1$.
Definition 2.5.
  1. (1)
    An ungraded representation $(V,\gamma)$ and $(V^{\prime},\gamma^{\prime})$ of $Cl_{n}$ are equivalent if there is a linear isomorphism $\phi\colon V\to V^{\prime}$ that preserves the actions of $Cl_{n}$.
  2. (2)
    A $\mathbb{Z}_{2}$-graded representation $(\hat{V},\hat{\gamma})$ and $(\hat{V}^{\prime},\hat{\gamma}^{\prime})$ of $Cl_{n}$ are equivalent if there is a linear isomorphism $\hat{\phi}\colon\hat{V}\to\hat{V}^{\prime}$ of degree 0 that preserves the actions of $Cl_{n}$.
Let $(\hat{V}=V^{0}\oplus V^{1},\,\hat{\gamma}=(\hat{\gamma}_{1},\ldots,\hat{\gamma}_{n}))$ be a $\mathbb{Z}_{2}$-graded representation of $Cl_{n}$ for a natural number $n\geq 2$. We can then consider $V^{0}=V^{1}=V$, and assume that the involution $\hat{\epsilon}$ and $\hat{\gamma}_{n}$ are of the following form:
$\hat{\epsilon}$ $=\begin{pmatrix}1_{V}&0\\ 0&-1_{V}\end{pmatrix},$ $\hat{\gamma}_{n}=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix}.$
Then, each $\hat{\gamma}_{i}$ has the following form:
$\hat{\gamma}_{i}=\begin{pmatrix}0&-\gamma_{i}\\ \gamma_{i}&0\end{pmatrix}$
for $i=1,\ldots,n-1$.
The linear maps $\gamma_{1},...,\gamma_{n}$ obtained here give a representation $V$ of $Cl_{n-1}$. However, it is obtained by specifying the concrete form of $\gamma_{n}$ and depends on its choice.
Proposition 2.6.
Let $\nu_{n}$ be the number of equivalence classes of the irreducible representations of $Cl_{n}$, and $\hat{\nu}_{n}$ that of the graded irreducible representations. Then,
$\nu_{n}$ =$=\begin{cases}1,&(n=0\mod 2)\\ 2,&(n=1\mod 2)\end{cases}$ $\hat{\nu}_{n}=\begin{cases}2,&(n=0\mod 2)\\ 1.&(n=1\mod 2)\end{cases}$
When there are two equivalence classes of irreducible representations, their dimensions are equal. Furthermore, if the dimension of the irreducible representation of $Cl_{n}$ is $d_{n}$ and the dimension of the graded irreducible representation is $\hat{d}_{n}$, then the following holds:
$d_{1}$ $=1,$ $d_{2}$ $=2,$ $d_{n+2}$ $=2d_{n},$
$\hat{d}_{1}$ $=2,$ $\hat{d}_{2}$ $=2,$ $\hat{d}_{n+2}$ $=2\hat{d}_{n}.$
$n$ $1$ $2$ $3$ $4$ $5$ $6$cdots $2k$ $12k+$ 1
$d_{n}$ 1 2 2 4 4 8 $\cdots$ $2^{k}$ $2^{k}$
$\hat{d}_{n}$ 2 2 4 4 8 8 $\cdots$ $2^{k}$ $2^{k+1}$
Proof.
See [2]. ∎
Definition 2.7.
Let $(\hat{V},\hat{\gamma})$ and $(\hat{V^{\prime}},\hat{\gamma}^{\prime})$ be $\mathbb{Z}_{2}$-graded representations of $Cl_{n}$ and $Cl_{m}$, respectively. Furthermore, let $\epsilon$ and $\epsilon^{\prime}$ be the involutions that define the $\mathbb{Z}_{2}$-gradings of $\hat{V}$ and $\hat{V}^{\prime}$, respectively. We then define the $\mathbb{Z}_{2}$-graded tensor product representation of the two $\mathbb{Z}_{2}$-graded representations $(\hat{V},\hat{\gamma})$ and $(\hat{V^{\prime}},\hat{\gamma}^{\prime})$ by the following pair.
$(\hat{V}\otimes\hat{V^{\prime}},((\hat{\gamma_{1}}\otimes 1_{\hat{V^{\prime}}},...,\hat{\gamma_{n}}\otimes 1_{\hat{V^{\prime}}}),(1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{1},...,1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{m})).$
Here, the $\mathbb{Z}_{2}$-grading of $\hat{V}\otimes\hat{V}^{\prime}$ is given by $\epsilon\otimes\epsilon^{\prime}$. Additionally, the product is defined as follows for $i,j=1,...,n$ and $k,l=1,...,m$.
$\hat{\gamma}_{i}\otimes 1_{\hat{V^{\prime}}}\circ\hat{\gamma}_{j}\otimes 1_{\hat{V^{\prime}}}=\hat{\gamma_{i}}\circ\hat{\gamma_{j}}\otimes 1_{\hat{V^{\prime}}},$
$\hat{\gamma_{i}}\otimes 1_{\hat{V^{\prime}}}\circ 1_{\hat{V}}\otimes\hat{\gamma^{\prime}_{k}}=\hat{\gamma_{i}}\otimes\hat{\gamma^{\prime}}_{k},$
$1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{k}\circ\hat{\gamma}_{i}\otimes 1_{\hat{V^{\prime}}}=-\hat{\gamma}_{i}\otimes\hat{\gamma^{\prime}}_{k},$
$1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{k}\circ 1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{l}=1_{\hat{V}}\otimes\hat{\gamma^{\prime}}_{k}\circ\hat{\gamma^{\prime}}_{l}.$
Proposition 2.8.
The following holds for the $\mathbb{Z}_{2}$-graded tensor product representation.
  1. (1)
    The isomorphism $Cl_{n}\otimes Cl_{m}\cong Cl_{n+m}$ holds. Furthermore, if $\hat{V}$ and $\hat{V}^{\prime}$ are $\mathbb{Z}_{2}$-graded representations of $Cl_{n}$ and $Cl_{m}$, respectively, then the $\mathbb{Z}_{2}$-graded tensor product representation $\hat{V}\otimes\hat{V}^{\prime}$ becomes a $\mathbb{Z}_{2}$-graded representation of $Cl_{n+m}$.
  2. (2)
    Suppose that $\hat{V}$ and $\hat{V}^{\prime}$ are $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{2n}$ and $Cl_{2m}$, respectively. Then, the $\mathbb{Z}_{2}$-graded tensor product representation $\hat{V}\otimes\hat{V}^{\prime}$ becomes a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n+2m}$
Proof.
For $(1)$, see [2] and [24]. $(2)$ is described in [13].
Here we illustrate examples of irreducible representations that will be used in the following chapters.
When $n=1$, let $V_{\pm}=\mathbb{C}$ and $Cl_{1}$ act by $\gamma_{1}=\pm 1$. These are two mutually inequivalent irreducible representations of $Cl_{1}$.
When $n=2$, one of the $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{2}$ can be given by considering the following action on $\hat{V}=\mathbb{C}\oplus\mathbb{C}$:
$\hat{\epsilon}$ $=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ $\hat{\gamma_{1}}$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ $\hat{\gamma_{2}}$ $=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix}.$
When $n=3$, one of the irreducible representations of $Cl_{3}$ is given by the following Pauli matrices:
$\gamma_{1}$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ $\gamma_{2}$ $=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix},$ $\gamma_{3}$ $=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
3. Formulation of topological $K$-theory via Fredholm operators
3.1. Definition of the $K$-group
In this section, we summarize the formulation of $K$-theory by using Fredholm operators, based on [3]. For other formulations of topological $K$-theory, see, for example, [1] and [20].
Definition 3.1.
For $n\geq 0$, let ^$\hat{\mathcal{H}}_{n}$ be a separable infinite-dimensional Hilbert space with the following properties:
  1. (1)
    It has $\mathbb{Z}_{2}$-grading and contains all $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{n}$ with infinite multiplicity. When $n=0$, the parts of degree 00 and $1$ are infinite-dimensional.
  2. (2)
    The above $\mathbb{Z}_{2}$-graded representation of $Cl_{n}$ can be extended to a $\mathbb{Z}_{2}$-graded representation of $Cl_{n+1}$.
Remark 3.2.
  1. (1)
    For any $n\geq 0$, ^$\hat{\mathcal{H}}_{n}$ exists. In fact, $Cl_{n+1}\otimes l^{2}$ is a separable infinite-dimensional Hilbert space that satisfies the properties of ^$\hat{\mathcal{H}}_{n}$
  2. (2)
    ^$\hat{\mathcal{H}}_{n}$ is isomorphic to each other by a bounded operator which preserves the action of $Cl_{n+1}$ and the inner product, and the whole such isomorphisms form a contractible space with respect to the topology given by the operator norm. (This is a consequence of Kuiper’s theorem [23].) Thus, ^$\hat{\mathcal{H}}_{n}$ is uniquely defined up to isomorphism.
Definition 3.3.
Let $Fred_{n}=Fred_{n}(\hat{\mathcal{H}}_{n})$ be the set of Fredholm operators $A\colon\hat{\mathcal{H}}_{n}\to\hat{\mathcal{H}}_{n}$ that are self-adjoint, degree 1, and anticommute with the actions $\gamma_{1},\gamma_{2},\ldots,\gamma_{n}$ of $Cl_{n}$. Then, we topologize $Fred_{n}$ by the operator norm.
Lemma 3.4.
For any $n\geq 0$, the following map is a homeomorphism:
$\phi\colon Fred_{n}(\hat{\mathcal{H}}_{n})\approx Fred_{n+2}(\hat{\mathcal{H}}\otimes\hat{S}),$ $\hat{A}\mapsto\hat{A}\otimes 1.$
Here, $\hat{S}$ is one of the $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{2}$.
Proof.
As the $\mathbb{Z}_{2}$-graded irreducible representation $\hat{S}$, we use the previously given example. Namely, $\hat{S}=\mathbb{C}\oplus\mathbb{C}$, and the grading $\hat{\epsilon}$ and $\hat{\gamma}$ are given as follows:
$\hat{\epsilon}$ $=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ $\hat{\gamma_{1}}$ $=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},$ $\hat{\gamma_{2}}$ $=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix}.$
It is clear from the definition that $\phi$ is continuous and injective. We will prove surjectivity. From the definition of the degree of ^$\hat{\mathcal{H}}_{n}\otimes\hat{S}$, the following is obtained:
(^$(\hat{\mathcal{H}}_{n}\otimes\hat{S})^{0}=\mathcal{H}_{n}^{0}\otimes S^{0}\oplus\mathcal{H}_{n}^{1}\otimes S^{1},$ (^$(\hat{\mathcal{H}}_{n}\otimes\hat{S})^{1}=\mathcal{H}_{n}^{0}\otimes S^{1}\oplus\mathcal{H}_{n}^{1}\otimes S^{0}.$
Let us arbitrarily take $\hat{B}\in Fred_{n+2}(\hat{\mathcal{H}}_{n}\otimes\hat{S})$ and we define $e_{0}={}^{t}\!(1,0)$ and $e_{1}={}^{t}\!(0,1)\in\mathbb{C}^{2}$. For any $a_{0}\in\mathcal{H}_{n}^{0}$, we describe the images of $a_{0}\otimes e_{0}$ and $a_{0}\otimes e_{1}$ under $\hat{B}$ as follows:
$\hat{B}(a_{0}\otimes e_{0})=x\otimes e_{0}+y\otimes e_{1},$ $\hat{B}(a_{0}\otimes e_{1})=z\otimes e_{0}+w\otimes e_{1}.$
We consider the following as a bilinear mapping $B_{1}$ on $\mathcal{H}_{n}^{0}\times S^{0}$:
$B_{1}\colon\mathcal{H}_{n}^{0}\times S^{0}\to\hat{\mathcal{H}}\otimes\hat{S},$ $(x,y)\mapsto\hat{B}(x\otimes y).$
Since $\hat{B}$ anticommutes with $1\otimes\hat{\gamma}_{1}$ and $1\otimes\hat{\gamma}_{2}$, the following holds:
00 0 $=(B_{1}\circ(1\times\hat{\gamma}_{1})+(1\otimes\hat{\gamma}_{1})\circ B_{1})(a_{0},e_{0})=B_{1}(a_{0},\hat{\gamma}_{1}e_{1})+1\otimes\hat{\gamma}_{1}(x\otimes e_{0}+y\otimes e_{1})$
$=(z\otimes e_{0}+w\otimes e_{1})-(x\otimes e_{1})+y\otimes e_{1}=(z+y)\otimes e_{0}+(w-x)\otimes e_{1}.$
From a similar calculation, we obtain the following:
$0=(B_{1}\circ(1\times\hat{\gamma}_{1})+(1\otimes\hat{\gamma}_{1})\circ B_{1})(a_{0},e_{0})=\sqrt{-1}(z-y)\otimes e_{0}+\sqrt{-1}(w-x)\otimes e_{1}.$
From the above calculation, it follows that $z=y=0$ and $w=x$, and thus, for some linear mapping $A_{1}\colon\mathcal{H}_{n}^{0}\to\mathcal{H}_{n}^{1}$, we can write $B_{1}=A_{1}\times 1$. Thus, from the universality of the tensor product, we obtain $\hat{B}|_{\mathcal{H}_{n}^{0}\otimes S^{0}}=A_{1}\otimes 1$.
By repeating exactly the same argument on $\mathcal{H}_{n}^{1}\otimes S^{1}$, $\mathcal{H}_{n}^{0}\otimes S^{1}$, and $\mathcal{H}_{n}^{1}\otimes S^{1}$, it is shown that $\hat{B}=\hat{A}\otimes 1$ for some self-adjoint linear mapping $\hat{A}\colon\hat{\mathcal{H}}\to\hat{\mathcal{H}}$ of degree $1$.
From the action of the Clifford algebra on ^$\hat{\mathcal{H}}_{n}$ and the anticommutativity of $\hat{B}$, it follows that $\hat{A}$ anticommutes with the action of $Cl_{n}$. Furthermore, from the Fredholm property of $\hat{B}$, it follows that $\hat{A}$ is Fredholm. That is, $\hat{A}\in Fred_{n}(\mathcal{H}_{n})$ and $\phi$ is surjective. Since $\phi$ is bijective and continuous, it follows from the open mapping theorem that $\phi$ is a homeomorphism. ∎
From Lemma 3.4, it follows that $Fred_{n}\approx Fred_{n+2}$.
When $n=0$, the direct summands of ^$\hat{\mathcal{H}}_{0}=\mathcal{H}_{0}^{0}\oplus\mathcal{H}_{0}^{1}$ are both infinite-dimensional separable Hilbert spaces. Therefore, they can be identified. That is, $\mathcal{H}_{0}=\mathcal{H}_{1}=\mathcal{H}$. In this case, the involution $\hat{\epsilon}$ and $\hat{A}\in Fred_{0}(\hat{\mathcal{H}})$ can be expressed as follows:
$\hat{\epsilon}=\begin{pmatrix}1_{\mathcal{H}}&0\\ 0&-1_{\mathcal{H}}\end{pmatrix},$ $\hat{A}=\begin{pmatrix}0&A^{*}\\ A&0\end{pmatrix}.$
Here, $A\colon\mathcal{H}\to\mathcal{H}$ is a Fredholm operator. By this correspondence, $Fred_{0}$ can be identified with the space of all Fredholm operators.
When $n=1$, the involution $\hat{\epsilon}$ and $\mathbb{Z}_{2}$-graded representation $\gamma_{1}$ of $Cl_{1}$ can be expressed as follows:
$\hat{\epsilon}=\begin{pmatrix}1_{\mathcal{H}}&0\\ 0&-1_{\mathcal{H}}\end{pmatrix},$ $\gamma_{1}=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix}.$
Then, $\hat{A}\in Fred_{1}(\hat{\mathcal{H}})$ has the following form:
$\hat{A}=\begin{pmatrix}0&A\\ A&0\end{pmatrix}.$
Here, $A\colon\mathcal{H}\to\mathcal{H}$ is a self-adjoint Fredholm operator. We define $Fred_{sa}(\mathcal{H})$ as the set of all self-adjoint Fredholm operators on $\mathcal{H}$, and topologize it with the relative topology induced from $Fred(\mathcal{H})$. By this correspondence, $Fred_{1}$ can be identified with $Fred_{sa}$.
Proposition 3.5.
The space $Fred_{sa}$ of self-adjoint Fredholm operators decomposes into the following three path-connected components:
  1. (1)
    The subspace $Fred_{sa}^{+}$, consisting of self-adjoint Fredholm operators that are positive in the Calkin algebra, forms a path-connected component of $Fred_{sa}$ and is contractible.
  2. (2)
    The subspace $Fred_{sa}^{-}$, consisting of self-adjoint Fredholm operators that are negative in the Calkin algebra, forms a path-connected component of $Fred_{sa}$ and is contractible.
  3. (3)
    The subspace $Fred_{sa}^{1}=Fred_{sa}\setminus(Fred_{sa}^{+}\cup Fred_{sa}^{-})$ forms a path-connected component of $Fred_{sa}$ and is not contractible.
Proof.
See [3]. ∎
Definition 3.6.
  1. (1)
    Let $\mathcal{F}_{1}(\mathcal{H}_{1})\subset Fred_{1}(\mathcal{H}_{1})$ be the subspace corresponding to $Fred_{sa}^{1}$.
  2. (2)
    Under the homeomorphism $Fred_{1}\approx Fred_{2k+1}$ given by Lemma 3.4, let $\mathcal{F}_{2k+1}$ be the subspace corresponding to $\mathcal{F}_{1}\subset Fred_{2k+1}$. Also, let $\mathcal{F}_{2k}=Fred_{2k}$.
  3. (3)
    For $n\geq 0$, let $\mathcal{F}_{n}^{*}\subset\mathcal{F}_{n}$ be the subspace of invertible operators.
Definition 3.7.
  1. (1)
    A pair $(X,Y)$ is defined as follows: $X$ is a topological space, and $Y$ is a closed subset of $X$. In particular, when $X$ is a compact Hausdorff space, the pair $(X,Y)$ is called a compact pair. Additionally, when $X$ is a CW-complex and $Y$ is a subcomplex, the pair $(X,Y)$ is referred to as a CW-pair.
  2. (2)
    Consider two pairs $(X,Y)$ and $(Z,W)$. A continuous map $f\colon X\to Z$ is called a map from the pair $(X,Y)$ to the pair $(Z,W)$ if the image of $Y$ under $f$ is contained in $W$. Such a map is denoted by $f\colon(X,Y)\to(Z,W)$.
  3. (3)
    Two maps $f,g\colon(X,Y)\to(Z,W)$ between the pairs $(X,Y)$ and $(Z,W)$ are homotopic if there exists a map $H\colon(X\times[0,1],Y\times[0,1])\to(Z,W)$ such that $H(x,0)=f(x)$ and $H(x,1)=g(x)$ for all $x\in X$. The quotient of the set of all maps from $(X,Y)$ to $(Z,W)$ under the homotopy relation is denoted by $[(X,Y),(Z,W)]$.
Definition 3.8.
Let $n\geq 0$ be a non-negative integer, and $(X,Y)$ a compact pair (or CW-pair). Then, $K^{-n}(X,Y)$ is defined as follows:
$K^{-n}(X,Y)=[(X,Y),(\mathcal{F}_{n},\mathcal{F}_{n}^{*})].$
In particular, when $Y=pt$, that is, for a pointed space X, the reduced $K$-group is defined as follows:
$\tilde{K}^{-n}(X)=K^{-n}(X,pt).$
It is known that the following holds:
$\tilde{K}^{-n}(X/Y)=K^{-n}(X,Y).$
The Abelian group structure on $K^{-n}(X,Y)$ is induced by the direct sum of the family of Fredholm operators by using an isomorphism ^$\hat{\mathcal{H}}_{n}\oplus\hat{\mathcal{H}}_{n}\cong\hat{\mathcal{H}}_{n}$
The operation of changing the $\mathbb{Z}_{2}$-grading on the Hilbert space ^$\hat{\mathcal{H}}_{n}$ is defined as replacing the involution $\varepsilon$, which determines the original $\mathbb{Z}_{2}$-grading, with its negative $-\varepsilon$. The resulting graded Hilbert space is denoted by $\Pi\hat{\mathcal{H}}_{n}$.
Based on this definition, the inverse of a family of Fredholm operators $\hat{A}\colon X\to\mathcal{F}_{n}$ in the $K$-group is given by
$\Pi\hat{A}\colon X\to\mathcal{F}_{n}(\Pi\hat{\mathcal{H}}_{n}),\quad\Pi\hat{A}_{x}=\hat{A}_{x}.$
That is, while the operator $\hat{A}$ itself remains unchanged, its inverse is defined through the change in the $\mathbb{Z}_{2}$-grading of the underlying Hilbert space.
By Lemma 3.4, there is a natural isomorphism $K^{-n}(X,Y)\cong K^{-n-2}(X,Y)$ of Abelian groups. By using this periodicity, we define $K^{n}(X,Y)$ and $\tilde{K}^{n}(X)$ for natural numbers $n\in\mathbb{N}$ as well.
The $K$-theory product $K^{-n}(X,Y)\times K^{-m}(X,Z)\to$ $K^{n-m}(X,Y\cup Z)$ is induced by
[…]
It is known that $K$-theory constitutes a generalized cohomology theory. Below, we recall the axioms of generalized cohomology.
Proposition 3.9.
The following statements hold:
  • Homotopy Axiom: Let $(X,Y)$ and $(Z,W)$ be compact pairs. If two maps $A,B:(X,Y)\to(Z,W)$ are homotopic, denoted $A\simeq B$, then the induced maps $A^{*},B^{*}:K^{*}(Z,W)\to K^{*}(X,Y)$ on the cohomology groups are equal.
  • Exactness Axiom: For any compact pair $(X,A)$, the following sequence
[…]
    is exact. Here, $i\colon A\to X$ and $j\colon X\to(X,A)$ are inclusion maps.
  • Excision Axiom: Let $(X,A)$ and $(X,B)$ be compact pairs. Then, for any $n\in\mathbb{Z}$, there is an isomorphism
    $i^{*}\colon K^{n}(A\cup B,B)\cong K^{n}(A,A\cap B),$
    where $i\colon(A,A\cap B)\to(A\cup B,B)$ is the inclusion map.
Proof.
The Homotopy Axiom follows directly from its definition. The Excision Axiom is a consequence of Kuiper’s Theorem [23]. The exactness axiom in the category of finite CW-complexes is derived from the suspension isomorphism and cofiber sequence. (The details on the suspension isomorphism will be discussed in the next section.) For details, see, for example, [6], [32]. For compact pairs, this can be understood by using the formulation in terms of vector bundles in topological $K$-theory [1]. ∎
3.2. Description of the generator of $K^{n}(D^{n},\partial D^{n})$
In this section, we let $I=[-\pi/2,\pi/2]$ and $\partial I=\{\pm\pi/2\}$.
When $\hat{A}:(X,Y)\to(\mathcal{F}_{n}(\mathcal{H}_{n+1}),\mathcal{F}_{n}^{*}(\mathcal{H}_{n+1}))$ is given, we define
$\mathrm{AS}(\hat{A}):(X\times I,Y\times\cup X\times\partial I)\to(\mathcal{F}_{n}(\mathcal{H}_{n+1}),\mathcal{F}^{*}_{n}(\mathcal{H}_{n+1}))$
as follows:
$\mathrm{AS}(\hat{A})(x,t)=\hat{A}_{x}\cos{t}-\hat{\gamma}_{n+1}\sin{t}.$
In their work [3], Atiyah and Singer establish an index theory for skew-adjoint Fredholm operators, derive the Bott periodicity, and relate the operator index to $K$-theory invariants. Building on this framework, we have:
Theorem 3.10.
$\mathrm{AS}(\hat{A}):(X\times I,Y\times I\cup X\times\partial I)\to(\mathcal{F}_{n}(\mathcal{H}_{n+1}),\mathcal{F}^{*}_{n}(\mathcal{H}_{n+1}))$ induces an isomorphism
$K^{-(n+1)}(X,Y)\cong K^{-n}(X\times I,Y\times I\cup X\times\partial I).$
Furthermore, the following isomorphism is also induced from AS:
$\tilde{K}^{-n-1}(X)\cong\tilde{K}^{-n}(SX)$
where $S$ is the suspension of a topological space, defined as $SX=(X\times I)/(pt\times I\cup X\times\partial I)$. These isomorphisms are called suspension isomorphisms.
Proof.
See [3]. ∎
Let $n$ be $1$ or $-1$. We define the operator $Sh_{n}$ on $l^{2}$ as follows.
$Sh_{n}((a_{1},a_{2},...))=\begin{cases}(0,a_{1},a_{2},...)&n=1,\\ (a_{2},a_{3},...)&n=-1.\end{cases}$
Clearly, the adjoint of $Sh_{-1}$ is $Sh_{1}$. Namely, $Sh_{-1}^{*}=Sh_{1}$. We define $\hat{l^{2}}$ as the direct sum of $l^{2}$, $\hat{l^{2}}=l^{2}\oplus l^{2}$.
Proposition 3.11.
The $K$-group of a point $pt$ is isomorphic to $\mathbb{Z}$, i.e., $K(pt)=[pt,\mathcal{F}_{0}(\hat{l}^{2})]\cong\mathbb{Z}$, and its generator is given as follows:
$\hat{Sh}=\begin{pmatrix}0&Sh_{1}\\ Sh_{-1}&0\end{pmatrix}.$
Proof.
The set $[pt,Fred(l^{2})]$ is bijective to the set $\mathbb{Z}$ of integers via the Fredholm index. Namely, $\text{Ind}:[pt,Fred(l^{2})]\cong\mathbb{Z}$. Since $\text{Ind}(Sh_{-1})=\dim{(\ker{(Sh_{-1}}))}-\dim{(\text{coker}(Sh_{-1}))}=1-0=1$, it follows that $Sh_{-1}$ provides a generator. In particular, from the correspondence between $Fred(l^{2})$ and $\mathcal{F}_{0}(l^{2})$, the operator $\hat{Sh}$ gives a generator for $K(pt)$. ∎
We define $l^{2}_{\geq 2}$ as the set of all elements in $l^{2}$ whose first component is 00. The space $l^{2}_{\geq 2}$ is a closed subspace, and the decomposition $l^{2}=\ker{Sh_{-1}}\oplus l^{2}_{\geq 2}$ holds. Note also that $Sh_{-1}$ induces an isomorphism $l^{2}_{\geq 2}\cong l^{2}$. The inverse mapping is given by $Sh_{1}:l^{2}\to l^{2}_{\geq 2}$. In particular, $\hat{Sh}:l^{2}_{\geq 2}\oplus l^{2}\cong l^{2}_{\geq 2}\oplus l^{2}$ is an invertible operator of degree $1$.
Let $A$, $B$, and $C$ be vector spaces. Suppose $f:A\to C$ and $g:B\to C$ are linear maps. Then the linear map $f+g:A\oplus B\to C$ is defined by
$(f+g)(a,b)=f(a)+g(b).$
Proposition 3.12.
Let $n\in\mathbb{N}$ be a natural number, and let $\hat{S}$ denote a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n}$. Then, with the zero map on $\hat{S}$ denoted by $\hat{0}$, the operator $\hat{0}+\hat{\gamma}$ on $\hat{S}\oplus\hat{\mathcal{H}_{2n}}$ provides a generator for $K^{-2n}(pt)=[pt,\mathcal{F}_{2n}(\hat{S}\oplus\hat{\mathcal{H}}_{2n})]$. Here, $\hat{\gamma}:\hat{\mathcal{H}_{2n}}\to\hat{S}\oplus\hat{\mathcal{H}_{2n}}$ is an invertible operator that anticommutes with the action of $Cl_{2n}$.
Proof.
We provide a proof by an induction on $n\in\mathbb{N}$.
  1. (1)
    Consider the case when $n=1$.
    Proposition 3.11 and Lemma 3.4 show that $\hat{Sh}\otimes 1\in[pt,\mathcal{F}_{2n}(l^{2}\otimes\hat{S})]$ provides a generator. Under the identification given by the isomorphisms $\mathbb{C}\otimes\hat{S}\cong\hat{S}$ and $Sh_{-1}\otimes 1\colon l^{2}_{\geq 2}\otimes\hat{S}\cong l^{2}\otimes\hat{S}$, which preserve the action of the Clifford algebra $Cl_{2}$, the operator $\hat{Sh}\otimes 1$ corresponds to $\hat{0}+\hat{\phi}\colon\hat{S}\oplus(\hat{l^{2}}\otimes\hat{S})\to\hat{S}\oplus(\hat{l^{2}}\otimes\hat{S})$. Here, $\hat{\phi}:\hat{l^{2}}\otimes\hat{S}\cong\hat{S}\oplus(\hat{l^{2}}\otimes\hat{S})$ is an invertible operator of degree $1$ induced by $\hat{Sh}\times 1:(l^{2}_{\geq 2}\oplus l^{2})\times\hat{S}\to(l^{2}_{\geq 2}\oplus l^{2})\times\hat{S}$. By construction, $\hat{\phi}$ anticommutes with the action of $Cl_{2}$. Therefore, the claim follows in the case $n=1$.
  2. (2)
    Let $n\in\mathbb{N}$ be a natural number greater than or equal to $2$, and assume that the claim holds up to $n-1$.
    The claim follows immediately from the induction hypothesis, Proposition 2.8, and Lemma 3.4.
Let $\hat{S}$ be a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n}$. Then, we define $\hat{\mu}_{\hat{S}}:D^{n}\to\text{End}(\hat{S})$ as follows.
$\hat{\mu}_{\hat{S}}(x_{1},...,x_{n})=\sum_{i=1}^{n}x_{i}\hat{\gamma}_{i}.$
Theorem 3.13.
Let $n\in\mathbb{Z}$ be a non-negative integer. There are the following isomorphisms:
$K^{-2n}(D^{2n},\partial D^{2n})\cong\mathbb{Z},$ $K^{-2n-1}(D^{2n+1},\partial D^{2n+1})\cong\mathbb{Z}.$
Let $\hat{S}$ be a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n}$. Then,
$\hat{\mu}_{\hat{S}}+\hat{\gamma}_{2n+1}=(\sum_{i=1}^{2n}x_{i}\hat{\gamma}_{i})+\hat{\gamma}_{2n+1}$
provides a generator for $K^{2n}(D^{2n},\partial D^{2n})\cong\mathbb{Z}$. Here, $\hat{\gamma}_{2n+1}\colon\hat{\mathcal{H}}_{2n}\to\hat{S}\oplus\hat{\mathcal{H}}_{2n}$ is an invertible operator that anticommutes with the action of $Cl_{2n}$.
Similarly, let $\hat{S}$ be a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n+1}$. Then,
$\hat{\mu}_{\hat{S}}+\hat{\gamma}_{2n+2}=(\sum_{i=1}^{2n+1}x_{i}\hat{\gamma}_{i})+\hat{\gamma}_{2n+2}$
provides a generator for $K^{-2n-1}(D^{2n+1},\partial D^{2n+1})\cong\mathbb{Z}$. Here, $\hat{\gamma}_{2n+2}\colon\hat{\mathcal{H}}_{2n+1}\to\hat{S}\oplus\hat{\mathcal{H}}_{2n+1}$ is an invertible operator that anticommutes with the action of $Cl_{2n+1}$.
Proof.
Since the arguments for even and odd degrees are similar, we provide a proof only for the even case. Given that $K^{-2n}(pt)\cong\mathbb{Z}$, applying the suspension isomorphism $n$ times yields $K(I^{n},\partial I^{n})\cong\mathbb{Z}$. Below, we explicitly compute $AS^{n}$. Let $k\in\mathbb{N}$ be a natural number. For computational convenience, we define $AS:K^{-k}(I^{k},\partial I^{k})\to K^{-k-1}(I^{k+1},\partial I^{k+1})$ by
$AS(\hat{A})(t_{1},...,t_{k+1})=\hat{A}_{(t_{1},...,t_{k})}\cos{t_{k+1}}+\hat{\gamma}_{k+2}\sin{t_{k+1}},$ $(t_{1},...,t_{k+1})\in[-\pi/2,\pi/2]^{k+1}.$
(As elements of the $K$-group, they differ at most by a sign multiple.)
Let $\hat{A}\in\mathcal{F}_{2n}(\hat{\mathcal{H}}_{2n})$. By the definition of the map $AS$, the expression for $AS^{2n}(\hat{A})$ is given by
$AS^{2n}(\hat{A})(t_{1},\dots,t_{2n})=\big{(}\cdots\big{(}(\hat{A}\cos{t_{1}}+\hat{\gamma}_{1}\sin{t_{1}})\cos{t_{2}}+\hat{\gamma}_{2}\sin{t_{2}}\big{)}\cdots\big{)}\cos{t_{2n}}+\hat{\gamma}_{2n}\sin{t_{2n}}.$
Here, for $j=0,...,2n,$ we define $x_{j}=x_{j}(t_{1},...,t_{2n})$ as follows:
$x_{0}=\cos{t_{1}}cos{t_{2}}\cdots\cos{t_{2n}},$
$x_{1}=\sin{t_{1}}\cos{t_{2}}\cos{t_{3}}\cdots\cos{t_{2n}},$
$\vdots$
$x_{2n-1}=\sin{t_{2n-1}}\cos{t_{2n}},$
$x_{2n}=\sin{t_{2n}}.$
This is the polar coordinate of the $n$-dimensional sphere $S^{2n}\subset\mathbb{R}^{2n+1}$, and the map
$I^{2n}\to\{(x_{0},...,x_{2n})\in\mathbb{R}^{2n+1}|x_{0}\geq 0\},$ $(t_{1},...,t_{2n})\mapsto(x_{0},,,,x_{2n})$
is a homeomorphism. Furthermore, the following map is also a homeomorphism
$\{(x_{0},...,x_{2n})|x_{0}\geq 0\}\to D^{2n},$ $(x_{0},...,x_{2n})\mapsto(x_{1},...,x_{2n}).$
By using these homeomorphisms, we identify $AS^{2n}(\hat{A})$ with the map
$AS^{2n}(\hat{A})\colon(D^{2n},\partial D^{2n})\to(\mathcal{F}_{0}(\hat{H}_{2n}),\mathcal{F}_{0}^{*}(\hat{H}_{2n})).$
This map has the following representation.
$AS^{2n}(\hat{A})(x_{1},...,x_{2n})=\sqrt{(1-\sum_{i=1}^{2n}{x_{i}}^{2}})\hat{A}+\sum_{i=1}^{2n}x_{i}\hat{\gamma}_{i}.$
If we let $\hat{S}$ be a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2n}$, then Proposition 3.12 shows that $[\hat{0}+\hat{\gamma}_{2n+1}]\in K^{-2n}(pt)=[pt,\hat{S}\oplus\mathcal{H}_{2n}]$ is a generator. Therefore,
$AS^{2n}(\hat{0}+\hat{\gamma}_{2n+1})=\hat{\mu}_{\hat{S}}+AS^{2n}(\hat{\gamma}_{2n+1})$
is given, and $AS^{2n}(\hat{\gamma}_{2n+1})\colon\hat{\mathcal{H}}_{2n}\to\hat{S}\oplus\hat{\mathcal{H}}_{2n}$ is an invertible operator that anticommutes with the action of $Cl_{2n}.$
3.3. Chern character and Push-forward
In this section, we define the Chern character for families of Fredholm operators. For the standard definition of the Chern character for vector bundles, see [19]. The definition of the even-degree Chern character $\operatorname{Ch}_{\text{even}}$ in terms of differential forms is given in [33], while the definition of the odd-degree Chern character $\operatorname{Ch}_{\text{odd}}$ is provided in [30]. Additionally, in this section, we define the push-forward map along the inclusion map from a point, which is necessary for the proof of the main theorem. For a more general definition of the push-forward map, see [20]. For the suspension isomorphism of singular cohomology, see, for example, [17].
For $n,m\in\mathbb{N}$ with $n\leq m$, we define the Grassmann manifold $G_{n}(\mathbb{C}^{m})$ as the set of all $n$-dimensional complex linear subspaces of $\mathbb{C}^{m}$. We define $BU(n)=G_{n}(\mathbb{C}^{\infty})$ by:
$BU(n)=G_{n}(\mathbb{C}^{\infty})=\bigcup_{m=n}^{\infty}G_{n}(\mathbb{C}^{m}).$
The inclusions $\mathbb{C}^{n}\subset\mathbb{C}^{n+1}\subset\dots$ yield inclusions $G_{m}(\mathbb{C}^{n})\subset G_{m}(\mathbb{C}^{n+1})\subset\dots$ and one can equip $G_{m}(\mathbb{C}^{\infty})$ with the direct limit topology.
The maps $G_{m}(\mathbb{C}^{n})\hookrightarrow G_{m+1}(\mathbb{C}^{n+1})$, which consist of adding the subspace generated by the last vector $e_{n+1}=(0,\dots,0,1)$, induce maps between the Grassmann manifolds $BU(m)\hookrightarrow BU(m+1)$. Let
$BU=\bigcup_{m=0}^{\infty}BU(m)$
with the inductive limit topology.
Theorem 3.14.
  1. (1)
    The integer coefficient cohomology of BU is given as follows:
    $H^{*}(BU;\mathbb{Z})\cong\mathbb{Z}[c_{1},c_{2},...],$ $c_{i}\in H^{2i}(BU;\mathbb{Z}).$
  2. (2)
    $\mathcal{F}_{0}$ and ×$\mathbb{Z}\times BU$ are homotopy equivalent:
    $\mathcal{F}_{0}\simeq\mathbb{Z}\times BU.$
Proof.
For the first claim, see [25] and [19]. ×$\mathbb{Z}\times BU$ and $\mathcal{F}_{0}$ are $\Omega$-spectra representing K-cohomology in the category of CW-complexes. Since an $\Omega$-spectrum representing a generalized cohomology $h^{*}$ on the category of CW-complexes is uniquely determined up to homotopy equivalence (((([6], [32])))), the second claim follows. ∎
Let $n\in\mathbb{N}$ be a natural number. Set $p_{0}=n$, and for each $k\in\mathbb{Z}$, we define $p_{k}\in H^{2k}(BU;\mathbb{Z})$ inductively by the following formula.
$p_{k}=c_{1}p_{k-1}-c_{2}p_{k-2}+\cdots+(-1)^{k}c_{k-1}p_{1}+(-1)^{k+1}kc_{k}.$
Furthermore, let us define the map $\phi:S\mathcal{F}_{0}\to\mathcal{F}_{1}$ as the mapping induced by the function $\mathcal{F}_{0}\times I\to\mathcal{F}_{1}$ given by $(A,t)\mapsto\hat{A}\cos t-\hat{\gamma}_{2}\sin t$. Since AS provides a suspension isomorphism, it follows that $\phi$ is a homotopy equivalence.
Definition 3.15.
  1. (1)
    For each $n\in\mathbb{N}$, we define the $n$-th universal Chern character $\operatorname{ch}_{n}^{\text{univ}}\in H^{2n}(\mathcal{F}_{0};\mathbb{Q})$ as follows:
    $\operatorname{ch}_{n}^{\text{univ}}=\frac{p_{n}}{n!}\in H^{2n}(\mathcal{F}_{0};\mathbb{Q}).$
  2. (2)
    For each $n\in\mathbb{N}$, we define $\operatorname{ch}_{n+\frac{1}{2}}^{\text{univ}}$ by the following formula:
    $\operatorname{ch}_{n+\frac{1}{2}}^{univ}=(\phi^{*})^{-1}(\delta(\operatorname{ch}_{n}^{univ}))\in H^{2n+1}(\mathcal{F}_{1};\mathbb{Q}),$
    Where $\delta\colon\tilde{H}^{*}(\mathcal{F}_{1})\cong\tilde{H}^{*+1}(S\mathcal{F}_{1})$ is the suspension isomorphism in singular cohomology.
Definition 3.16.
Let $n\in\mathbb{Z}$ be a non-negative integer.
  1. (1)
    For any $[\hat{A}]\in K(X)$, we define $\operatorname{ch}_{n}([\hat{A}])\in H^{2n}(X;\mathbb{Q})$ as follows:
    $\operatorname{ch}_{n}([\hat{A}])=\hat{A}^{*}\operatorname{ch}_{n}^{univ}\in H^{2n}(X;\mathbb{Q}),$ $n\geq 1,$
    $\operatorname{ch}_{0}([\hat{A}])=Ind([A_{x}])\in H^{0}(X;\mathbb{Q}),$ $x\in X.$
  2. (2)
    For any $[\hat{A}]\in K^{1}(X)$, we define $\operatorname{ch}_{n+\frac{1}{2}}([\hat{A}])\in H^{2n+1}(X;\mathbb{Q})$ as follows:
    $\operatorname{ch}_{n+\frac{1}{2}}([\hat{A}])=\hat{A}^{*}\operatorname{ch}_{n+\frac{1}{2}}^{univ}\in H^{2n+1}(X;\mathbb{Q}).$
  3. (3)
    The homomorphisms $\operatorname{Ch}_{even}$ and $\operatorname{Ch}_{odd}$ from the $K$-group to singular cohomology are defined as follows:
    $\operatorname{Ch}_{even}\colon K(X)\to H^{even}(X),$ $\operatorname{Ch}_{even}([A])=\operatorname{ch}_{0}([A])+\operatorname{ch}_{1}([A])+\cdots,$
    $\operatorname{Ch}_{odd}\colon K^{1}(X)\to H^{odd}(X),$ $\operatorname{Ch}_{odd}([A])=\operatorname{ch}_{\frac{1}{2}}([A])+\operatorname{ch}_{\frac{3}{2}}([A])+\cdots.$
$\operatorname{Ch}_{even}$ and $\operatorname{Ch}_{odd}$ are called the Chern characters.
Let $X$ be a pointed space. The inclusion map from the basepoint to $X$ is denoted by $i$. From the definition of the Chern character, it follows that $i^{*}\operatorname{Ch}(\hat{A})=\operatorname{Ch}(i^{*}A)$. Therefore, the Chern character induces a homomorphism from the reduced $K$-group to the reduced singular cohomology:
$\operatorname{Ch}_{even}\colon\tilde{K}(X)\to\tilde{H}^{even}(X),$ $\operatorname{Ch}_{odd}\colon\tilde{K}^{1}(X)\to\tilde{H}^{odd}(X).$
Proposition 3.17.
The suspension isomorphism and the Chern character are commutative. That is, the following diagram commutes:
$\tilde{K}^{0}(X)$$\tilde{H}^{\mathrm{even}}(X;\mathbb{Q})$$\tilde{K}^{1}(SX)$$\tilde{H}^{\mathrm{odd}}(SX;\mathbb{Q})$$\scriptstyle\operatorname{Ch}_{\mathrm{even}}$$\scriptstyle\delta$$\scriptstyle AS$$\scriptstyle\operatorname{Ch}_{\mathrm{odd}}$
Proof.
Given a map $\hat{A}:X\to\mathcal{F}_{0}$, we define $S\hat{A}\colon SX\to S\mathcal{F}_{0}$ as the map induced by the following mapping:
$X\times I\to\mathcal{F}_{1},$ $(x,t)\mapsto(\hat{A}_{x},t).$
Notice that, by definition, the equality $\phi\circ S\hat{A}=AS(\hat{A})$ holds. Additionally, the suspension isomorphism in singular cohomology is essentially a connecting homomorphism. Therefore, for each integer $n\in\mathbb{Z}$, the claim follows from the calculations below.
$\delta(\operatorname{ch}_{n}(\hat{A}))=\delta(\hat{A}^{*}\operatorname{ch}_{n}^{univ})=(SA)^{*}\delta(\operatorname{ch}_{n}^{univ})=(SA)^{*}\phi^{*}(\operatorname{ch}_{n+\frac{1}{2}}^{univ})$
$=AS(\hat{A})^{*}\operatorname{ch}_{n+\frac{1}{2}}^{univ}=\operatorname{ch}_{n+\frac{1}{2}}(AS(\hat{A})).$
Definition 3.18.
Let $X$ be a compact Hausdorff space. Suppose there exists a closed neighborhood of some $x\in X$ that is homeomorphic to the $n$-dimensional closed disk $D^{n}$ (this closed neighborhood is also denoted by $D^{n}$). Then, $i_{*}^{k}\colon K^{0}({x})\to K^{n}(X)$ is defined from the following diagram.
[…]
Here, $j:(D^{n},\partial D^{n})\to(X,X\setminus\text{int}(D^{n}))$ and $k:(X,X\setminus\text{int}(D^{n}))\to(X,X\setminus\{x\})$ are inclusion maps. We call $i_{*}^{k}$ the push-forward.
Proposition 3.19.
Let $X$ be a compact Hausdorff space. Suppose there exists a closed neighborhood of some $x\in X$ that is homeomorphic to the $n$-dimensional closed disk $D^{n}$. Then, the following diagram is commutative.
[…]
Here, $i_{*}^{H}$ is the push-forward in singular cohomology.
Proof.
This follows from the construction of the push-forward and Proposition 3.17. ∎
4. Expression of Chern character by using Fermi points
4.1. Approximation of family of Fredholm operators
In this section, we introduce a finite-dimensional approximation of a Fredholm operator following [15].
Let $X$ be a topological space, $\hat{\mathcal{H}}$ a $\mathbb{Z}_{2}$-graded separable infinite-dimensional Hilbert space, and $\hat{A}\colon X\to\mathcal{F}(\hat{\mathcal{H}})$ a continuous map with $x_{0}\in X$ and $0\in Spec(A_{x_{0}}).$
Proposition 4.1.
There exists a positive real number $\mu>0$ and an open neighborhood U of $x_{0}$ that satisfy the following three properties:
  1. (1)
    $Spec(\hat{A}_{x})\cap[0,\mu)$ is a finite set.
  2. (2)
    For $\lambda\in Spec(\hat{A}_{x})\cap[0,\mu)$, the eigenspace corresponding to $\lambda$ is finite-dimensional.
  3. (3)
    For any $x\in U$, $\dim(\bigoplus_{|\lambda|<\mu}\ker(\hat{A}_{x_{0}}-\lambda))=\dim(\bigoplus_{\lvert\lambda\rvert<\mu}\ker(\hat{A}_{x}-\lambda)).$
Proof.
Because $\hat{A}_{x_{0}}$ is a Fredholm operator, $0\in Spec(\hat{A}_{x_{0}})$ is a discrete spectrum, so $[0,\mu)\cap Spec(\hat{A}_{x_{0}})$ is a finite set and there exists a positive real number $\mu>0$ such that any $\lambda\in Spec(\hat{A}_{x_{0}})\cap[0,\mu)$ has finite multiplicity. Therefore, due to the continuity of the eigenvalues, we can obtain an open neighborhood of $x_{0}$ that fulfills the assertion of the proposition. ∎
In what follows, we assume that $U$ and $\mu>0$ are given by the above Proposition 4.1.
Definition 4.2.
  1. (1)
[…]
  2. (2)
    We define the family of vector spaces (^$(\hat{\mathcal{H}},\hat{A})$ over $U$ as the set whose vector space at each point $x$ is (^$(\hat{\mathcal{H}},\hat{A}_{x})$:
[…]
  3. (3)
    We define $F$ as the set of all $x\in X$ such that $\hat{A}_{x}$ has 00 in its spectrum:
    $F=\{x\in X\mid 0\in\mathrm{Spec}(\hat{A}_{x})\}.$
    We call this $F$ the Fermi set with respect to $\hat{A}$. Similarly, for an ungraded Fredholm operator, we define the Fermi set in the same manner.
The purpose of this subsection is to establish:
Proposition 4.3.
(^$(\hat{\mathcal{H}},\hat{A})_{<\mu}\subset U\times\hat{\mathcal{H}}$ is a $\mathbb{Z}_{2}$-graded complex vector bundle of finite rank.
Proof.
Fix $x\in U$ and we will show that the family of vector spaces (^$(\hat{\mathcal{H}},\hat{A})_{<\mu}$ can be locally trivialized in some open neighborhood of $x$.
For each $y\in U$, we define the projection $\pi_{y}:\hat{\mathcal{H}}\to(\hat{\mathcal{H}},\hat{A}_{y})_{<\mu}$. Let $i:(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}\hookrightarrow\hat{\mathcal{H}}$ be an inclusion map. We have $(1_{\hat{\mathcal{H}}}-\pi_{x})\circ(1_{\hat{\mathcal{H}}}-\pi_{x})\circ i=1_{(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}^{\perp}}:(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}^{\perp}\to(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}^{\perp}$, and noting that the set of invertible bounded operators is an open set in $\mathcal{B}(\hat{\mathcal{H}})$, we obtain the following assertion.
For any $y\in V$, there exists an open neighborhood $V\subset U$ of $x$ such that $(1_{\hat{\mathcal{H}}}-\pi_{x})\circ(1_{\hat{\mathcal{H}}}-\pi_{y})\circ i:(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}^{\perp}\to(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}^{\perp}$ is a linear isomorphism.
Then, the continuous map $\overline{\pi_{x}}:(\hat{\mathcal{H}},\hat{A})_{<\mu}|_{V}\to V\times(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}$ induced by $\pi_{x}$ gives a local trivialization in some open neighborhood of $x$. In fact, for any $y\in V$, the following diagram commutes.
[…]
From the commutative diagram $\pi_{x}\circ i:\ker((1_{\hat{\mathcal{H}}}-\pi_{x})\circ(1_{\hat{\mathcal{H}}}-\pi_{y}))\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}$ is an isomorphism.
By definition, (^$(\hat{\mathcal{H}},\hat{A}_{y})_{<\mu}=\ker(1_{\hat{\mathcal{H}}}-\pi_{y})\subset\ker((1_{\hat{\mathcal{H}}}-\pi_{x})\circ(1_{\hat{\mathcal{H}}}-\pi_{y}))$ and $\dim((\hat{\mathcal{H}},\hat{A}_{x})_{<\mu})=\dim((\hat{\mathcal{H}},\hat{A}_{y})_{<\mu})<\infty$ hold. Therefore, $\ker((1_{\hat{\mathcal{H}}}-\pi_{x})\circ(1_{\hat{\mathcal{H}}}-\pi_{y}))=\ker(1_{\hat{\mathcal{H}}}-\pi_{y})$ holds.
Hence, $\overline{\pi_{x}}:(\hat{\mathcal{H}},\hat{A})_{<\mu}|_{V}\to V\times(\hat{\mathcal{H}},\hat{A}_{x})_{<\mu}$ is a local trivialization. ∎
Definition 4.4.
By Proposition 4.3, we define a finite-dimensional approximation $\hat{A}_{<\mu}:(\hat{\mathcal{H}},\hat{A})_{<\mu}\to(\hat{\mathcal{H}},\hat{A})_{<\mu}$ of $\hat{A}$ by setting $\hat{A}_{<\mu}(x)=\hat{A}(x)$. In an ungraded case, we similarly define $(\mathcal{H},A)_{<\mu}$ and $A_{<\mu}:(\mathcal{H},A)_{<\mu}\to(\mathcal{H},A)_{<\mu}$.
4.2. Sign coordinate
Definition 4.5.
[…]
  1. (1)
[…]
    Here, $\simeq$ means fiber homotopy equivalence.
  2. (2)
    When a sign coordinate $(V;a_{1},...,a_{2k})$ $(V\subset U)$ can be taken in $x\in F$, define the following sign.
    $\operatorname{sign}(x)=\frac{\frac{\partial(a_{1},...,a_{2k+1})}{\partial(x_{1},...,x_{2k})}}{\lvert\frac{\partial(a_{1},...,a_{2k})}{\partial(x_{1},...,x_{2k})}\rvert}\in\{-1,1\}.$
    Here $(V;x_{1},...,x_{2k})$ is an oriented coordinate of $X$, and $\frac{\partial(a_{1},...,a_{2k})}{\partial(x_{1},...,x_{2k})}$ is the Jacobian.
Definition 4.6.
Let $k$ be a non-negative integer. Let $X$ be a ($12k+$ 1)-dimensional compact, oriented, differentiable manifold, and let $A\colon X\to Fred_{sa}(\mathcal{H})$ be a continuous map. We fix $\mu>0$ and an open set $U\subset X$, as described in Proposition 4.1, such that $U$ contains $F$ and the space $(\mathcal{H},A_{x})_{<\mu}$ is defined for all $x\in U$. We assume that $\dim(\mathcal{H},A_{x})_{<\mu}=2^{k}$ for all $x\in F$. We also fix an action of $Cl_{2k}$ on $(\mathcal{H},A_{x})_{<\mu}$ such that it becomes an irreducible representation.
  1. (1)
    We say that a coordinate $(V;a_{1},...,a_{2k+1})$ $(x\in V\subset U)$ of $X$ is an (odd) sign coordinate of $x$ when $A_{<\mu}|_{V}\simeq\sum_{n=1}^{2k+1}a_{i}\gamma_{i}:(\mathcal{H},A)_{<\mu}|_{V}\to(\mathcal{H},A)_{<\mu}|_{V}.$
    Here, $\simeq$ means fiber homotopy equivalence.
  2. (2)
    When a sign coordinate $(V;a_{1},...,a_{2k+1})$ $(V\subset U)$ can be taken in $x\in F$, define the following sign.
    $\operatorname{sign}(x)=\frac{\frac{\partial(a_{1},...,a_{2k+1})}{\partial(x_{1},...,x_{2k+1})}}{\lvert\frac{\partial(a_{1},...,a_{2k+1})}{\partial(x_{1},...,x_{2k+1})}\rvert}\in\{-1,1\}.$
    Here $(V;x_{1},...,x_{2k+1})$ is an oriented coordinate of $X$, and $\frac{\partial(a_{1},...,a_{2k+1})}{\partial(x_{1},...,x_{2k+1})}$ is the Jacobian.
Definition 4.7.
An element $x\in F$ of the Fermi set is called a Fermi point if a sign coordinate can be defined for it.
Let $U_{1}$ and $U_{2}$ be open neighborhoods of the origin in $\mathbb{R}^{n}$. Consider a diffeomorphism $f\colon U_{1}\to U_{2}$ satisfying $f(0)=0$ and $\det{T}>0$, where $T=Jf(0)$ is the Jacobian matrix of $f$ at the origin. The image of $T$, denoted $Im(T)$, is the image of the linear map $T$. We define $f_{1}\colon Im(T)\to U_{2}$ by $f_{1}=f\circ T^{-1}$, so that $f=f_{1}\circ T$. Here, $Jf$ and $Jf_{1}$ denotes the Jacobian matrix of $f$ and $f_{1}$, respectively.
Lemma 4.8.
$Jf_{1}(0)=E_{n}$, where $E_{n}$ is the $n\times n$ identity matrix.
Proof.
This is clear from the definition of $f_{1}$ and the differentiation of composite and inverse functions. ∎
Lemma 4.9.
  1. (1)
    Let $U_{1},U_{2}\subset\mathbb{R}^{n}$ be open neighborhoods of 0, $f:U_{1}\to U_{2}$ a smooth map such that $f(0)=0$ and $Jf(0)=E_{n}$. Then $f|_{D(r)}:D(r)\to D(r^{\prime})$ is a diffeomorphism and there exists a positive real numbers $r,r^{\prime}>0\,(r^{\prime}\geq r>0)$ such that $f|_{D(r)}\simeq i$. Here, $D(r)$ is an open ball of radius $r$ centered at the origin, and $i:D(r)\to D(r^{\prime})$ is the inclusion map.
  2. (2)
    Let $T\in SL(n;\mathbb{R})$ be a linear transformation. Then for any $r>0$, there exists $0<r^{\prime\prime}<r$ such that
    $\operatorname{Im}(T|_{D(r^{\prime\prime})})\subset D(r),$
    and the map $T\colon D(r^{\prime\prime})\to D(r)$ is homotopic to the inclusion $i\colon D(r^{\prime\prime})\hookrightarrow D(r)$; that is, $T\simeq i$.
Proof.
  1. (1)
    For any real number $t\in\mathbb{R}$, we define a homotopy $f_{t}\colon U_{1}\to\mathbb{R}^{n}$ by the formula
    $f_{t}(x)=tx+(1-t)f(x).$
    This definition ensures that $f_{t}(0)=0$ and $f_{t}(1)=1$.
    Next, we define a map $F\colon U_{1}\times[0,1]\to\mathbb{R}^{n}\times[0,1]$ by $F(x,t)=(f_{t}(x),t)$. By definition, the image of the point $(0,t)$ under the map $F$ is $F(0,t)=(0,t)$. Furthermore, the Jacobian matrix of $F$ at $(0,t)$ has the form
    $JF(0,t)=\begin{pmatrix}E_{n}&*\\ 0&1\end{pmatrix},$
    where the asterisk represents some entries whose explicit form is irrelevant in this context. Since the Jacobian matrix is non-singular, the inverse function theorem guarantees the existence of positive real numbers $r_{t}>0$ and $\delta_{t}>0$ such that the restricted map
    $F|_{D(r_{t})\times[t-\delta_{t},t+\delta_{t}]}\colon D(r_{t})\times[t-\delta_{t},t+\delta_{t}]\to\mathbb{R}^{n}\times[0,1]$
    is a diffeomorphism onto its image. Due to the compactness of $[0,1]$, there exists a monotonely increasing sequence $t_{1},...,t_{n}\in[0,1]$ such that $\bigcup_{i=1}^{n}[t_{i}-\delta_{t_{i}},t_{i}+\delta_{t_{i}}]=[0,1]$. If we set $r=\min\{r_{t_{1}},\dots,r_{t_{n}}\}$, then $F|_{D(r)\times[0,1]}:D(r)\times[0,1]\to\mathbb{R}^{n}\times[0,1]$ is a diffeomorphism onto its image. The subspace $F(D(r)\times[0,1])\subset\mathbb{R}^{n}\times[0,1]$ is compact, so there exists a positive real number $r^{\prime}>0$ such that $F(D(r)\times[0,1])\subset D(r^{\prime})\times[0,1]$. ($0<r\leq r^{\prime}$ since $F(D(r)\times\{1\})=D(r)\times\{1\}\subset D(r^{\prime})\times\{1\}$.) Therefore, $F$ is a homotopy that gives $f|_{D(r)}\simeq i$.
  2. (2)
    Since $SL(n;\mathbb{R})$ is path-connected, the path connecting $T$ and $E_{n}$ gives the homotopy.
Given a $\mathbb{Z}_{2}$-graded representation $\hat{S}$ of $Cl_{2k}$, let $\hat{\mu}_{S}\colon D^{2k}\to End(\hat{S})$ be defined as
$\hat{\mu}_{S}(x_{1},...,x_{2k+1})=\sum_{i=1}^{2k+1}x_{i}\hat{\gamma}_{i}.$
The orthogonal complement of the embedding $\hat{S}\subset\mathcal{H}_{2k}$ as a graded representation is denoted as $\hat{S}^{\perp}$, and the action of $Cl_{2k+1}$, which is an extension of the action of $Cl_{2k}$, on $\hat{S}^{\perp}\cong\mathcal{H}_{2k}$ is denoted as $\hat{\gamma}^{\perp}$.
Theorem 4.10.
Suppose that a continuous mapping
[…]
satisfies the following assumptions:
  1. (1)
    The Hilbert space $\hat{\mathcal{H}}$ admits a direct sum decomposition ^$\hat{\mathcal{H}}=\hat{S}_{+}\oplus\hat{S}_{+}^{\perp}$, and accordingly, the operator $\hat{A}(x)$ can be decomposed as
    $\hat{A}(x)=\hat{a}(x)+\hat{a}^{\perp}(x),$
    where $\hat{a}(x)\colon\hat{S}_{+}\to\hat{S}_{+}\subset\hat{\mathcal{H}}$ and $\hat{a}^{\perp}(x)\colon\hat{S}_{+}^{\perp}\to\hat{\mathcal{H}}$.
  2. (2)
    For any $x\in D^{2k}$, we have $\hat{a}^{\perp}(x)\in\mathcal{F}^{*}_{0}(\mathcal{\hat{\mathcal{H}}})$.
  3. (3)
    There exists a smooth mapping $f=(f_{1},...,f_{2k})\colon D^{2k}\to\mathbb{R}^{2k}$ such that:
    • For any $x\in D^{2k}$, the map $\hat{a}$ admits the expression $\hat{a}(x)=\sum_{i=1}^{2k}f_{i}(x)\hat{\gamma}_{i}$.
    • $f^{-1}(\{0\})=\{0\}$.
    • $\det Jf(0)>0$.
In this case, under an isomorphism $K^{0}(D^{2k},\partial D^{2k})\cong\mathbb{Z}$, the map $\hat{A}$ represents the generator $1\in\mathbb{Z}$.
Proof.
  1. (1)
    Consider the case where $f$ satisfies $\det{Jf(0)}=1$.
    Under the third assumption, Lemma 4.9 guarantees the existence of positive real numbers $r^{\prime}\geq r>0$ such that the map $f:D(r)\to D(r^{\prime})$ is a diffeomorphism and is homotopy equivalent to the inclusion $i:D(r)\to D(r^{\prime})$.
    If necessary, $r$ and $r^{\prime}$ may be replaced so as to be $r\leq r^{\prime}\leq 1$. From the long exact sequence in cohomology for the triple $(D(r^{\prime}),D(r^{\prime})\setminus\operatorname{int}(D(r)),\partial D(r^{\prime}))$, we obtain the following isomorphism:
    $j^{*}\colon K^{-1}(D(r^{\prime}),D(r^{\prime})\setminus\text{int}(D(r))\cong K^{-1}(D(r^{\prime}),\partial D(r^{\prime})).$
    Therefore, from this isomorphism and the excision theorem, we obtain the following isomorphism:
    $f^{*}=i^{*}\colon K^{-1}(D(r^{\prime}),\partial D(r^{\prime}))\cong K^{-1}(D(r),\partial D(r)).$
    Then we obtain the following sequence of equalities:
    $i^{*}([\hat{\mu}_{S}+\hat{\gamma}^{\perp}|_{D(r^{\prime})}])$ $=f^{*}([\hat{\mu}_{S}+\hat{\gamma}^{\perp}])=[\hat{\mu}_{S}\circ f+\hat{\gamma}^{\perp}\circ f]$
    $=\left[(\sum_{i=1}^{2k+1}f_{i}\hat{\gamma}_{i})+\hat{a}^{\perp}\right]=i^{*}([\hat{A}|_{D(r^{\prime})}]).$
    Thus, it follows that $[\hat{\mu}_{S}+\hat{\gamma}^{\perp}|_{D(r^{\prime})}]=[\hat{A}|_{D(r^{\prime})}]$.
    We note that $D^{2k+1}=D(1)$. From the long exact sequence in cohomology for the triple $(D^{2k+1},D^{2k+1}\setminus\operatorname{int}(D(r^{\prime})),\partial D^{2k+1})$, we obtain the following isomorphism:
    $j^{*}\colon K^{-1}(D^{2k+1},D^{2k+1}\setminus\operatorname{int}(D(r^{\prime})))\cong K^{-1}(D^{2k+1},\partial D^{2k+1}).$
    Under the assumption that $f^{-1}(\{0\})=\{0\}$, the operator $\hat{A}$ is invertible on $D^{2k+1}\setminus\operatorname{int}(D(r^{\prime})$. Therefore, $\hat{A}$ represents an element in $K^{-1}(D^{2k+1},D^{2k+1}\setminus\operatorname{int}(D(r^{\prime})))$. From this isomorphism and the excision theorem, we obtain the following isomorphism:
    $i^{*}\colon K^{-1}(D^{2k+1},\partial D^{2k+1})\cong K^{-1}(D(r^{\prime}),\partial D(r^{\prime})).$
    Consequently, from $i^{*}([\hat{\mu}_{S}+\hat{\gamma}^{\perp}])=i^{*}([\hat{A}])$, we deduce that $[\hat{\mu}_{S}+\hat{\gamma}^{\perp}]=[\hat{A}]$.
  2. (2)
    Consider the case of a general $f$.
    From Lemma 4.8, we can decompose $f$ as $f=f_{1}\circ T^{-1}$, where $T$ is a homeomorphism which maps the origin to itself. Lemma 4.9 guarantees the existence of positive real numbers $r,r^{\prime}$ with $0<r\leq r^{\prime}$ such that the map
    $f_{1}:D(r)\to D(r^{\prime})$
    is a diffeomorphism which is homotopic to the inclusion $i$, that is, $f\simeq i$. By Lemma 4.9, There exists a positive real number $r^{\prime\prime}>0$ such that
    $D(r^{\prime\prime})\subset\operatorname{Im}(T|_{D(r)})$
    and the map
    $T^{-1}:D(r^{\prime\prime})\to D(r)$
    is homotopic to the inclusion $i:D(r^{\prime\prime})\to D(r)$. Consequently, we have
    $f=f_{1}\circ T^{-1}:D(r^{\prime\prime})\to D(r^{\prime}).$
    By Lemma 4.8, this implies that
    $f^{*}=f_{1}^{*}:K^{-1}(D(r^{\prime}),\partial D(r^{\prime}))\to K^{-1}(D(r^{\prime\prime}),\partial D(r^{\prime\prime})).$
    Hence, the claim follows from the discussion in (1).
Let $\hat{S}_{+}$ and $\hat{S}_{-}$ be two $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{2k+2}$ that are inequivalent to each other. (If necessary, the $\mathbb{Z}_{2}$-graded irreducible representations of $Cl_{2k+1}$ obtained by ignoring the action of $\gamma_{2k+2}\in Cl_{2k+2}$ are assumed to be identical.) Given a $\mathbb{Z}_{2}$-graded representation $\hat{S}$ of $Cl_{2k+2}$, let $\hat{\mu}_{S}\colon D^{2k+1}\to End(\hat{S})$ be defined as
$\hat{\mu}_{S}(x_{1},...,x_{2k+1})=\sum_{i=1}^{2k+1}x_{i}\hat{\gamma}_{i}.$
The orthogonal complement of the embedding $\hat{S}\subset\mathcal{H}_{2k+2}$ as a graded representation is denoted as $\hat{S}^{\perp}$, and the action of $Cl_{2k+3}$, which is an extension of the action of $Cl_{2k+2}$, on $\hat{S}^{\perp}\cong\mathcal{H}_{2k+2}$ is denoted as $\hat{\gamma}^{\perp}$.
Theorem 4.11.
Suppose that a continuous mapping
[…]
satisfies the following assumptions:
  1. (1)
    The Hilbert space $\hat{\mathcal{H}}$ admits an orthogonal decomposition ^$\hat{\mathcal{H}}=\hat{S}_{+}\oplus\hat{S}_{+}^{\perp}$, and accordingly, the operator $\hat{A}(x)$ can be decomposed as a sum
    $\hat{A}(x)=\hat{a}(x)+\hat{a}^{\perp}(x),$
    where $\hat{a}(x)\colon\hat{S}_{+}\to\hat{S}_{+}\subset\hat{\mathcal{H}}$ and $\hat{a}^{\perp}(x)\colon\hat{S}_{+}^{\perp}\to\hat{\mathcal{H}}$.
  2. (2)
    For any $x\in D^{2k+1}$, we have $\hat{a}^{\perp}(x)\in\mathcal{F}^{*}_{1}(\mathcal{\mathcal{H}})$.
  3. (3)
    There exists a smooth mapping $f=(f_{1},...,f_{2k+1})\colon D^{2k+1}\to\mathbb{R}^{2k+1}$ such that:
    • For any $x\in D^{2k+1}$, the map $\hat{a}$ admits the expression $\hat{a}(x)=\sum_{i=1}^{2k+1}f_{i}(x)\hat{\gamma}_{i}$.
    • $f^{-1}(\{0\})=\{0\}$.
    • $\det Jf(0)>0$.
In this case, under an isomorphism $K^{-1}(D^{2k+1},\partial D^{2k+1})\cong\mathbb{Z}$, the map $\hat{A}$ represents the generator $1\in\mathbb{Z}$.
Proof.
The proof proceeds similarly in the even-dimensional case. ∎
Remark 4.12.
Consider the $\mathbb{Z}_{2}$-graded representation $\hat{S}$ of the Clifford algebra $Cl_{2k+1}$, defined as $\hat{S}=S^{0}\oplus S^{1}$, where $S^{0}=S^{1}$. Under this representation, we express the following matrices:
$\hat{\epsilon}$ $=\begin{pmatrix}1_{S}&0\\ 0&-1_{S}\end{pmatrix},$ $\hat{\gamma}_{i}$ $=\begin{pmatrix}0&\gamma_{i}\\ \gamma_{i}&0\end{pmatrix},$ $\hat{\gamma}_{2k+2}$ $=\begin{pmatrix}0&-\sqrt{-1}\cdot 1_{S}\\ \sqrt{-1}\cdot 1_{S}&0\end{pmatrix}.$
where $i$ ranges from 1 to $12k+$ 1. Here, $\gamma_{1},\dots,\gamma_{2k+1}$ are self-adjoint homomorphisms on $S$, which collectively provide an ungraded representation of $Cl_{2k+1}$.
In this framework, we define the operators $\hat{a}(x)$ and $\hat{\mu}(x)$ as follows:
$\hat{a}(x)$ $=\begin{pmatrix}0&\sum_{i=1}^{2k+1}f_{i}(x)\gamma_{i}\\ \sum_{i=1}^{2k+1}f_{i}(x)\gamma_{i}&0\end{pmatrix},$ $\hat{\mu}(x)$ $=\begin{pmatrix}0&\sum_{i=1}^{2k+1}x_{i}\gamma_{i}\\ \sum_{i=1}^{2k+1}x_{i}\gamma_{i}&0\end{pmatrix}.$
From these definitions, it follows that a statement concerning $\hat{a}(x)$ can be equivalently reformulated as a statement about $a(x)$.
We define $Fred_{sa}^{1,*}(\mathcal{H})\subset Fred_{sa}^{1}(\mathcal{H})$ as the subspace of invertible operators.
Let $S_{+}$ and $S_{-}$ be two ungraded irreducible representations of $Cl_{2k+1}$ that are inequivalent to each other. Furthermore, consider a separable infinite-dimensional Hilbert space $\mathcal{H}$, and take the direct sum $S_{+}\oplus\mathcal{H}$ with $S_{+}$. In this case, the following corollary is obtained.
Corollary 4.13.
Consider a continuous map
$A\colon(D^{2k+1},\partial D^{2k+1})\to(Fred_{sa}^{1}(S_{+}\oplus\mathcal{H}),Fred_{sa}^{1,*}(S_{+}\oplus\mathcal{H})).$
We denote the $S_{+}$-component and the $\mathcal{H}$-component of $A$ by $a$ and $a^{\perp}$, respectively. That is, we write
$A(x)=a(x)+a^{\perp}(x),$
where $a(x)\colon S_{+}\to S_{+}\subset\mathcal{H}$ and $a^{\perp}(x)\colon S_{+}^{\perp}\to\mathcal{H}$. We assume that $A$ satisfies the following conditions.
  1. (1)
    For any $x\in D^{2k+1}$, we have $a^{\perp}(x)\in Fred_{sa}^{1,*}(S_{+}\oplus\mathcal{H})$.
  2. (2)
    There exists a smooth mapping $f=(f_{1},...,f_{2k+1})\colon D^{2k+1}\to\mathbb{R}^{2k+1}$ such that:
    • For any $x\in D^{2k+1}$, the map a admits the expression $a(x)=\sum_{i=1}^{2k+1}f_{i}(x)\gamma_{i}$.
    • $f^{-1}(\{0\})=\{0\}$.
    • $\det Jf(0)>0$.
In this case, under an isomorphism $K^{-1}(D^{2k+1},\partial D^{2k+1})\cong\mathbb{Z}$, the map $A$ represents the generator $1\in\mathbb{Z}$.
Proof.
The claim follows from the correspondence between $Fred_{sa}^{1}$ and $\mathcal{F}_{1}$, along with Theorem 4.11 and Remark 4.12. ∎
4.3. Proof of the main theorem
Let $X$ be a compact, oriented, differentiable manifold of dimension $12k+$ 1. Let  $A\colon X\to Fred_{sa}^{1}(\mathcal{H})$ be a continuous map and $F$ the Fermi set of $A$. Let $(V,b=(b_{1},\dots,b_{2k+1}))$ be an oriented coordinate of $X$, where $V$ is a coordinate neighborhood of $x\in F$ and $b:V\to\mathbb{R}^{2k+1}$ satisfies $b(x)=(0,\dots,0)$. If necessary, by suitably reselecting $V$, we can assume that the image $b(V)$ under $b$ coincides with the disk $D^{2k+1}$.
Through this $b$, we identify $b^{*}\colon K^{-1}(D^{2k+1},\partial D^{2k+1})\cong K^{-1}(V,\partial V).$ Under this identification, the generator $1$ corresponds to
$b^{*}[(\sum_{i=1}^{2k+1}x_{i}\gamma_{i})+\gamma^{\perp}]=[(\sum_{i=1}^{2k+1}b_{i}\gamma_{i})+\gamma^{\perp}].$
Lemma 4.14.
Let $X$ be a compact, oriented, differentiable manifold of dimension $12k+$ 1. Let  $A\colon X\to Fred_{sa}^{1}(\mathcal{H})$ be a continuous map and $F$ the Fermi set of $A$. Let $(V,a=(a_{1},...,a_{2k+1})$ be a sign coordinate of $A$ at $x\in F$. In this case, when $(V,b=(b_{1},...,b_{2k+1}))$ is an oriented coordinate,
$[(\sum_{i=1}^{2k+1}b_{i}\gamma_{i})+\gamma^{\perp}]=\operatorname{sign}(x)[(\sum_{i=1}^{2k+1}a_{i}\gamma_{i})+\gamma^{\perp}]$
holds.
Proof.
Let $f_{i}$ denote the projection of $a\circ b^{-1}$ onto the $i$-th component, for $i=1,\dots,2k+1$. By using $f_{i}$, we can express $a\circ b^{-1}$ as $(f_{1},\dots,f_{2k+1})$. Under this definition, the $K$-group classes satisfy the relation:
$[(\sum_{i=1}^{2k+1}x_{i}\gamma_{i})+\gamma^{\perp}]=\operatorname{sign}(x)[(\sum_{i=1}^{2k+1}f_{i}\gamma_{i})+\gamma^{\perp}].$
Consequently, we can derive the following sequence of equalities by applying the pullback $b^{*}$:
$[(\sum_{i=1}^{2k+1}b_{i}\gamma_{i})+\gamma^{\perp}]=b^{*}([(\sum_{i=1}^{2k+1}x_{i}\gamma_{i})+\gamma^{\perp}])$ $=\operatorname{sign}(x)b^{*}([(\sum_{i=1}^{2k+1}f_{i}\gamma_{i})+\gamma^{\perp}])$
$=\operatorname{sign}(x)[(\sum_{i=1}^{2k+1}a_{i}\gamma_{i})+\gamma^{\perp}].$
Lemma 4.15.
Let $X$ be a $(2k+1)$-dimensional compact, oriented differentiable manifold. Let  $A\colon X\to Fred_{sa}^{1}(\mathcal{H})$ be a continuous map and $F$ the Fermi set of $A$. We assume that the Fermi set $F=\{x\}$ is a singleton set. Then, the following holds.
  1. (1)
    The condition that a sign coordinate can be taken at $x$ is equivalent to the following: there exists a coordinate $(V;a=(a_{1},...,a_{2k+1}))$ around $x$ such that $f(V)=D^{2k+1}$, and the pullback $(a^{-1})^{*}([A|_{V}])\in K^{1}(D^{2k+1},\partial D^{2k+1})\cong\mathbb{Z}$ is a generator.
  2. (2)
    Suppose that a sign coordinate $(V;a=(a_{1},...,a_{2k+1}))$ can be taken around a point $x\in F$ on the manifold $X$.Under this condition, the following equality holds in the $K$-group:
    $i_{*}^{K}(1)=\operatorname{sign}(x)[A]\in K^{1}(X)$
    Here, $i:F\to X$ is an inclusion map from the Fermi set into $X$, and the element $1$ is identified with the generator of $K^{0}(F)$.
Proof.
  1. (1)
    Suppose there exists a coordinate $(V;a=(a_{1},...,a_{2k+1}))$ $(D^{2k+1}=f(V))$ around a point $x\in X$, such that $f(V)=D^{2k+1},$ and the $K$-group class $([A|_{D^{2k+1}}])\in K^{-1}(D^{2k+1},\partial D^{2k+1})\cong\mathbb{Z}$ is a generator. This assumption implies that $V$ is mapped onto the disk $D^{2k+1}$ via the coordinate function $f$, and $A$ restricted to this disk generates the $K$-group.
    If necessary, we may replace $a=(a_{1},...,a_{2k+1})$ with $(-a_{1},...,a_{2k+1})$ to adjust the orientation. Under this adjustment, the following equality holds:
    $(a^{-1})^{*}([A|_{V}])=[(\sum_{i=1}^{2k+1}x_{i}\gamma_{i})+\gamma^{\perp}]).$
    Consequently, it follows that: $[A|_{V}]=[(\sum_{i=1}^{2k+1}a_{i}\gamma_{i})+\gamma^{\perp}]).$ Here, the equality indicates that $A_{<\mu}|_{V}\simeq\sum_{i=1}^{2k+1}a_{i}\gamma_{i}$. The converse—that every sign coordinate satisfies these conditions—follows from Lemma 4.14.
  2. (2)
    Take an oriented coordinate $(V,b=(b_{1},...,b_{2k+1})$ $(f(V)=D^{2k+1})$ around $x$, and identify $V$ with $D^{2k+1}$. Then,
    $\tau(1)=[(\sum_{i=1}^{2k+1}b_{i}\gamma_{i})+\gamma^{\perp}])=\operatorname{sign}(x)[A|_{D^{2k+1}}].$
    Here, $\tau$ is the suspension isomorphism $\tau\colon K^{0}(F)\cong K^{1}(D^{2k+1},\partial D^{2k+1})$.
    Recall that the Fermi set $F$ is defined as the set of points $x\in X$ where the operator $A_{x}$ is singular, i.e., $F=\{x\in F|0\in Spec(A_{x})\}.$ By this definition, the $K$-group class $[A]$ can be regarded as an element of $K^{1}(X,X\setminus F)$, since $A$ is invertible on $X\setminus F$. With this understanding, the statement follows from the commutativity of the diagram below.
[…]
    where $j\colon(D^{n},\partial D^{n})\to(X,X\setminus\text{int}(D^{n}))$ and $k\colon(X,X\setminus\text{int}(D^{n}))\to(X,X\setminus\{x\})$ are an inclusion maps.
Lemma 4.16.
Let $X$ be a compact, oriented, differentiable manifold, and $\mathcal{H}$ a separable infinite-dimensional Hilbert space. Let $A\colon X\to\mathrm{Fred}_{sa}(\mathcal{H})$ be a continuous mapping. In this case, the points in $F$ that are Fermi points are isolated points.
Proof.
Suppose $x\in F$ has a sign coordinate but is not an isolated point. Then, for any sign coordinate of $x$, denoted $(U,x=(x_{1},\dots,x_{n}))$, there exists $x^{\prime}\in F\cap U\setminus\{x\}$. However, by the definition of a sign coordinate, the only point in $U$ where $A_{<\mu}=0$ holds is $x$. This implies that no such $x^{\prime}\neq x$ can exist in $U\cap F$, leading to a contradiction. ∎
Lemma 4.17.
Let $X$ be a compact, oriented, differentiable manifold, and $\mathcal{H}$ a separable infinite-dimensional Hilbert space. Let $A\colon X\to Fred_{sa}(\mathcal{H})$ be a continuous mapping, and let us assume that each point of the Fermi set $F$ is a Fermi point. Then, $F$ is a finite set.
Proof.
Since $A$ is a continuous map, $F$ is a closed set. Therefore, $V=X\backslash F$ is an open set. From the Lemma 4.16, for any $x\in F$, there exists an open neighborhood $U_{x}$ of $x$ such that $U_{x}\cap F=\{x\}$ holds. In this case, $X=V\cup\bigcup_{x\in F}U_{x}$ is an open covering of $X$, so due to the compactness, $X$ can be covered by a finite number of these. Then, $F\cap V=\phi$ means that $F$ is a finite set. ∎
Theorem 4.18 (Theorem 1.2).
Let $k$ be a non-negative integer. Let $X$ be a compact, oriented, $(2k+1)$-dimensional differentiable manifold, and  $A\colon X\to\mathcal{F}_{1}(\mathcal{H})$ a continuous map. Then, the following holds.
  1. (1)
    When each point $x\in F$ is a Fermi point, the sum of signs $\sum_{x\in F}\operatorname{sign}(x)$ depends only on $[A]\in K^{1}(X)$.
  2. (2)
    Under the assumption of $(1)$, the following equality holds.
    $\int_{X}\operatorname{ch}_{k+\frac{1}{2}}([A])=\sum_{x\in F}\operatorname{sign}(x)\in\mathbb{Z}.$
Proof.
The sum $\sum_{x\in F}\operatorname{sign}(x)$ is finite by Lemma 4.17. Since the integral of the Chern character depends only on the homotopy class of $A$, it is sufficient to prove the second claim. From Lemma 4.15, we can prove the case where $F=\{x\}$. From the Proposition 3.19, the following diagram commutes.
$K^{0}(F)$$H^{\mathrm{even}}(F;\mathbb{Q})$$K^{1}(X)$$H^{\mathrm{odd}}(X;\mathbb{Q})$$\scriptstyle\operatorname{Ch}$$\scriptstyle i_{*}^{H}$$\scriptstyle i_{*}^{K}$$\scriptstyle\operatorname{Ch}$
Therefore, we have
$\operatorname{ch}_{k+\frac{1}{2}}(i_{*}^{K}(1))=\operatorname{Ch}(i_{*}^{K}(1))=i_{*}^{H}(\operatorname{Ch}(1))=\mathrm{vol},$
where the volume form $\mathrm{vol}$ represents the fundamental top-degree form on the oriented compact manifold $X$, which integrates to $1$. Consequently, we obtain
$\int_{X}\operatorname{ch}_{k+\frac{1}{2}}(i_{*}^{K}(1))=1.$
By Lemma 4.15, it follows that
$\int_{X}\operatorname{ch}_{k+\frac{1}{2}}([A])=\operatorname{sign}(x).$
Similarly, the even-degree case follows.
Theorem 4.19 (Theorem 1.1).
[…]
  1. (1)
    When a sign coordinate is taken at each point $x\in F,$ the sum of signs $\sum_{x\in F}\operatorname{sign}(x)$ depends only on $[\hat{A}]\in K^{0}(X)$.
  2. (2)
    Under the assumption of $(1)$, the following equality holds.
    $\int_{X}\operatorname{ch}_{k}([\hat{A}])=\sum_{x\in F}\operatorname{sign}(x)\in\mathbb{Z}.$
4.4. Specific calculation example
Below, we present specific computational examples based on Theorem 4.19 and Theorem 4.18. In §§§4.4.1, we demonstrate that Theorem 4.18 corresponds to a generalization of the spectral flow. In this section, we denote $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$.
4.4.1. Example $1$
We define $(a,b)\colon S^{1}\to\mathbb{R}^{2}$ as follows:
$a(k)$ $=\frac{3}{2}+\cos k,$
$b(k)$ $=\sin k.$
We utilize the results of local models established in [14].
We define the following mappings by using the local model:
$\hat{H}$ $=\hat{H}_{\text{loc}}\circ(a,b)\times 1_{S^{1}}\colon T^{2}=S^{1}\times S^{1}\to\text{Herm}(\mathbb{C}^{2}),$
$\hat{H}^{\#}$ $=\hat{H}^{\#}_{\text{loc}}\circ(a,b)\colon S^{1}\to Fred_{\text{sa}}(l^{2}(\mathbb{N},\mathbb{C}^{2})).$
The Fermi set is $F=\{\pi\}$. We can approximate $\hat{H}^{\#}$ on $(-1,1)\times\{0\}\subset\mathbb{R}^{2}$ as follows:
$(\hat{H}^{\#}_{\text{loc}})_{<\sqrt{2}}=b.$
We have $\det Jb(\pi)=\cos\pi=-1<0$, where we define $J(b(k))=\frac{\partial b}{\partial k}(k)\in\mathbb{R}$. We conclude by the inverse function theorem that $b(k)$ serves as a sign coordinate of $S^{1}$ in a neighborhood of $\pi$. Thus, we obtain the following result by applying Theorem 4.18 with $\mathcal{H}=l^{2}(\mathbb{N},\mathbb{C}^{2})$:
$\int_{X}\operatorname{ch}_{\frac{1}{2}}([H^{\#}])=-1.$
We confirm that this result is consistent with $sf([H^{\#}])=-1$.
4.4.2. Example $2$
In §§§4.4.2 through §§§4.4.4, we use the local model $\hat{H}_{loc}\colon\mathbb{R}\times\mathbb{C}^{2}\times S^{1}\to\text{Herm}(\mathbb{C}^{4})$ given by the following equation:
$\hat{H}_{loc}(a,b,c,k)=\begin{pmatrix}a&\bar{c}-e^{ik}&0&\bar{b}\\ c-e^{-ik}&-a&\bar{b}&0\\ 0&b&a&-c+e^{-ik}\\ b&0&-\bar{c}+e^{ik}&-a\end{pmatrix}.$
We define $\hat{H}^{\#}_{loc}\colon\mathbb{R}\times\mathbb{C}^{2}\to Fred_{\text{sa}}(l^{2}(\mathbb{N},\mathbb{C}^{4}))$ as follows:
$\hat{H}^{\#}_{loc}(a,b,c)=H^{\#}_{(a,b,c)}\in Fred_{\text{sa}}(l^{2}(\mathbb{N},\mathbb{C}^{4})).$
We refer to §6 for the definition of $H^{\#}_{(a,b,c)}$. Below, we utilize the result of Proposition 6.8.
We note that, in the neighborhood of each point $\{0\}\times\{0\}\times D^{2}\subset\mathbb{R}\times\mathbb{C}^{2}$, the following holds:
$(\hat{H}^{\#}_{loc})_{<\sqrt{3}}=\text{Re}(b)\sigma_{1}+\text{Im}(b)\sigma_{2}+a\sigma_{3}.$
We let $k=(x,y+\sqrt{-1}z,w)\in\mathbb{R}\times\mathbb{C}\times\mathbb{R}$. We define $(a,b,c)\colon S^{3}\subset\mathbb{R}\times\mathbb{C}\times\mathbb{R}\to\mathbb{R}\times\mathbb{C}^{2}$ as follows:
$a(k)=x,$ $b(k)=y+\sqrt{-1}z$ , $c(k)=w-1.$
In this case, we use the local model to define $\hat{H}^{\#}=\hat{H}^{\#}_{loc}\circ(a,b,c)\colon S^{3}\to Fred_{\text{sa}}(l^{2}(\mathbb{N},\mathbb{C}^{4}))$.
By the spectral analysis of the local model, a point $k\in F$ in the Fermi set of $\hat{H}^{\#}$ satisfies:
$a(k)$ $=0,$ $b(k)=0$ , $\lvert c(k)\rvert$ $\leq 1.$
Therefore, we obtain $F=\{(0,0,1)\}$.
We define $J(\text{Re}(b(k)),\text{Im}(b(k)),a(k))$ as the Jacobian matrix. The determinant is as follows:
$\det J(\text{Re}(b(k)),\text{Im}(b(k)),a(k))=1.$
By applying the inverse function theorem, we establish that $(\text{Re}(b),\text{Im}(b),a)$ serves as a sign coordinate of $S^{3}$ in the neighborhood of $(0,0,1)$. Therefore, we apply the Theorem 4.18 with $\mathcal{H}=l^{2}(\mathbb{N},\mathbb{C}^{4})$ to obtain:
$\int_{S^{3}}\operatorname{ch}_{\frac{3}{2}}([H^{\#}])=1.$
4.4.3. Example $3$
We define $(a,b,c)\colon T^{3}=S^{1}\times S^{1}\times S^{1}\to\mathbb{R}\times\mathbb{C}^{2}$ as follows:
$a(k)$ $=-\cos(k_{2}+k_{3}),$
$b(k)$ $=(-\cos(k_{1})-\cos(k_{2}))+\sqrt{-1}(\sin(k_{1})+\sin(k_{2})),$
$c(k)$ $=-1-e^{-\sqrt{-1}k_{3}},$
where we denote $k=(k_{1},k_{2},k_{3})\in T^{3}$. In this case, we use the local model to define the following:
$\hat{H}$ $=\hat{H}_{loc}\circ(a,b,c)\times 1_{S^{1}}\colon T^{4}=T^{3}\times S^{1}\to\text{Herm}_{0}(\mathbb{C}^{4}),$
$\hat{H}^{\#}$ $=\hat{H}^{\#}_{loc}\circ(a,b,c)\colon T^{3}\to Fred_{\text{sa}}(l^{2}(\mathbb{N},\mathbb{C}^{4})).$
($\hat{H}$ is a Hamiltonian corresponding to a four-dimensional topological insulator in class AI [29],[31].)
We identify the Fermi set $F$ of $\hat{H}^{\#}$ by the conditions:
$a(k)=0,$ $b(k)=0$ , $\lvert c(k)\rvert$ $\leq 1.$
By the spectral analysis of the local model, a point $k\in F$ in the Fermi set of $\hat{H}^{\#}$ satisfies:
$F=\left\{\left(\frac{2\pi}{3},\frac{4\pi}{3},\frac{7\pi}{6}\right),\left(\frac{4\pi}{3},\frac{2\pi}{3},\frac{5\pi}{6}\right)\right\}.$
We let $J(\text{Re}(b(k)),\text{Im}(b(k)),a(k))$ denote the Jacobian matrix. Its determinant is
$\det J(\text{Re}(b(k)),\text{Im}(b(k)),a(k))=\sin(k_{1}-k_{2})\sin(k_{2}+k_{3}).$
For any $x\in F$, the determinant satisfies $\det J(x)=-\frac{\sqrt{3}}{2}<0$. Therefore, we apply the inverse function theorem and conclude that, in the neighborhood of each point in $F$, $(\text{Re}(b),\text{Im}(b),a)$ serves as a sign coordinates of $T^{3}$. Consequently, we apply Theorem 4.18 with $\mathcal{H}=l^{2}(\mathbb{N},\mathbb{C}^{4})$ and obtain:
$\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\#}])=-2.$
4.4.4. Example $4$
We define $\mathcal{H}=l^{2}(\mathbb{N},\mathbb{C}^{2})$ and ^$\hat{\mathcal{H}}=l^{2}(\mathbb{N},\mathbb{C}^{4})=l^{2}(\mathbb{N},\mathbb{C}^{2})\oplus l^{2}(\mathbb{N},\mathbb{C}^{2})=\mathcal{H}\oplus\mathcal{H}$, and we specify the involution as follows:
$\hat{\epsilon}=\begin{pmatrix}1_{\mathcal{H}}&0\\ 0&-1_{\mathcal{H}}\end{pmatrix}.$
We adopt the local model from the previous example with $a=0$. Specifically, we define $\hat{H}^{\#}_{loc}\colon\mathbb{C}^{2}\to Fred_{\text{sa}}(\hat{\mathcal{H}})$ as follows:
$\hat{H}^{\#}_{loc}(b,c)=H^{\#}_{(0,b,c)}\in Fred_{\text{sa}}(\hat{\mathcal{H}}).$
We observe that $\hat{H}^{\#}_{loc}$ anticommutes with $\hat{\epsilon}$, which implies that it is a map of degree $1$. Therefore, for each $k\in\mathbb{C}^{2}$, we have $\hat{H}^{\#}_{loc}(k)\in\mathcal{F}(\hat{\mathcal{H}})$.
The finite-dimensional approximation $(\hat{H}^{\#}_{\mathrm{loc}})_{<\sqrt{3}}$ of $\hat{H}^{\#}_{\mathrm{loc}}$, defined over a neighborhood of each point in $\{0\}\times D^{2}\subset\mathbb{C}^{2}$, acts on $\mathbb{C}^{2}$. The space $\mathbb{C}^{2}$ is a $\mathbb{Z}_{2}$-graded irreducible representation of $Cl_{2}$ via the action of the Pauli matrices $\sigma_{1}$ and $\sigma_{2}$. In this case, $(\hat{H}^{\#}_{\mathrm{loc}})_{<\sqrt{3}}$ has the following expression:
$(\hat{H}^{\#}_{loc})_{<\sqrt{3}}=\text{Re}(b)\sigma_{1}+\text{Im}(b)\sigma_{2}.$
We let $k=(y+\sqrt{-1}z,w)\in\mathbb{C}\times\mathbb{R}$. We define $(b,c)\colon S^{2}\subset\mathbb{C}\times\mathbb{R}\to\mathbb{C}^{2}$ as follows:
$b(k)=y+\sqrt{-1}z,$ $c(k)=w-1.$
In this case, we use the local model to define $\hat{H}^{\#}=\hat{H}^{\#}_{loc}\circ(b,c)\colon S^{2}\to\mathcal{F}(\hat{\mathcal{H}})$.
By the spectral analysis of the local model, a point $k\in F$ in the Fermi set of $\hat{H}^{\#}$ satisfies:
$b(k)=0,$ $\lvert c(k)\rvert$ $\leq 1.$
Thus, we obtain $F=\{(0,1)\}$.
We define $J(\text{Re}(b(k)),\text{Im}(b(k)))$ as the Jacobian matrix. We compute the determinant as follows:
$\det J(\text{Re}(b(k)),\text{Im}(b(k)))=1.$
By applying the inverse function theorem, we conclude that $(\text{Re}(b),\text{Im}(b))$ serves as a sign coordinate of $S^{2}$ in the neighborhood of $(0,1)$. Therefore, we apply Theorem 4.19 and obtain:
$\int_{S^{2}}\operatorname{ch}_{2}([\hat{H}^{\#}])=1.$
4.5. Application
Let $X$ be a compact, oriented, $3$-dimensional differentiable manifold. We suppose that a smooth involution $\tau$ acts on $X$ and assume that $\tau$ reverses the orientation of $X$.
We consider a continuous map $A\colon X\to Fred_{sa}^{1}(l^{2})$. We assume that a sign coordinate $(V;a_{1},a_{2},a_{3})$ can be defined at each isolated point of the Fermi set. We also assume that the vector space arising from the finite-dimensional approximation becomes the representation space of an irreducible representation of $Cl_{3}$ via the action of the Pauli matrices $\sigma_{1},\sigma_{2},\text{and }\sigma_{3}$. Furthermore, we assume that the finite-dimensional approximation $A_{<\mu}|_{V}$ of $A$ has the following expression:
$A_{<\mu}|_{V}=\sum_{i=1}^{3}a_{i}\gamma_{i}\colon(\mathcal{H},A)_{<\mu}|_{V}\to(\mathcal{H},A)_{<\mu}|_{V}.$
We define $K$ as the complex conjugation operator acting on $l^{2}$, specifically representing it as an operator that maps a vector $\psi=(a_{1},a_{2},...)\in l^{2}$ to its complex conjugate $\overline{\psi}$, i.e.,
$K(a_{1},a_{2},...)=(\bar{a_{1}},\bar{a_{2}},...).$
In the following discussion, we assume that $KA_{x}K=A_{\tau(x)}$ holds for any $x\in X$.
Lemma 4.20.
If we can take a sign coordinate $(U,x=(x_{1},x_{2},x_{3}))$ at $a\in F$, then $(\tau(U),\tau^{*}x=(\tau^{*}x_{1},-\tau^{*}x_{2},\tau^{*}x_{3}))$ is a sign coordinate at $\tau(a)\in F$.
Proof.
This is evident by using the assumption that $KA_{x}K=A_{\tau(x)}$ and $K\sigma_{1}K=\sigma_{1}$, $K\sigma_{2}K=-\sigma_{2}$, and $K\sigma_{3}K=-\sigma_{3}$. ∎
Lemma 4.21.
At fixed points of $\tau$ in the Fermi set $F$, no sign coordinates can be taken.
Proof.
Assume that $a\in F$ is a fixed point with a sign coordinate $(U,x=(x_{1},x_{2},x_{3}))$. If necessary, we can assume that $U=\tau(U)$ by taking $U$ sufficiently small. From the Lemma 4.20, $(\tau(U),\tau^{*}x=(\tau^{*}x_{1},-\tau^{*}x_{2},\tau^{*}x_{3}))$ is the sign coordinate at the fixed point $a$, and on $U$ we obtain the following formula:
$A_{<\mu}=x_{1}\sigma_{1}+x_{2}\sigma_{2}+x_{3}\sigma_{3}=\tau^{*}x_{1}\sigma_{1}-\tau^{*}x_{2}\sigma_{2}+\tau^{*}x_{3}\sigma_{3}.$
Therefore, on $U$, the formula $x_{1}=\tau^{*}x_{1}$ holds. Differentiating with respect to a, we get $(dx_{1})_{a}=(d\tau^{*}x_{1})_{a}=-(dx_{1})_{a}$. Therefore, we have $(dx_{1})_{a}=0$, which contradicts the fact that $(U,x=(x_{1},x_{2},x_{3}))$ forms a coordinate system. ∎
Proposition 4.22.
Under the above assumptions, we have $\int_{X}\operatorname{ch}_{\frac{3}{2}}([A])\in 2\mathbb{Z}$.
Proof.
From Lemma 4.21, there exists no fixed point with sign coordinates. If $a\in F$ has a sign coordinate, then by Lemma 4.20, $\tau(a)$ also has a sign coordinate. Thus the statement follows. ∎
5. Bulk-edge correspondence for class AI four-dimensional topological insulators
5.1. Real vector bundle
In this section, we summarize real vector bundles in the sense of Atiyah as presented in [4]. We explain this concept by following the approach of [11].
Definition 5.1.
We define an involution space to be a pair $(X,\tau)$ of a topological space $X$ and an involution $\tau$, namely a homeomorphism such that $\tau^{2}=1_{X}$.
We consider the Grassmann manifold $G_{m}(\mathbb{C}^{n})$. For each $m$-plane $\Sigma=\langle v_{1},\ldots,v_{m}\rangle$, we define its conjugate as $\overline{\Sigma}=\langle\overline{v}_{1},\ldots,\overline{v}_{m}\rangle$. We define an involution $\varrho:G_{m}(\mathbb{C}^{n})\to G_{m}(\mathbb{C}^{n})$ by $\varrho(\Sigma)=\overline{\Sigma}$. Furthermore, because the inclusions $G_{m}(\mathbb{C}^{n})\hookrightarrow G_{m}(\mathbb{C}^{n+1})$ commute with $\varrho$, the involution $\varrho$ extends to the infinite-dimensional Grassmann manifold $G_{m}(\mathbb{C}^{\infty}).$
Definition 5.2.
Let us consider a complex vector bundle $\pi\colon E\to X$ over an involution space $(X,\tau)$. We define a real vector bundle in the sense of Atiyah as a complex vector bundle equipped with a homeomorphism $\Theta\colon E\to E$ that satisfies the following conditions:
  1. (1)
    The projection is compatible with the involution, i.e., the equation
    $\pi\circ\Theta=\tau\circ\pi$
    holds.
  2. (2)
    The map $\Theta$ is anti-linear on each fiber, meaning that for any $\lambda\in\mathbb{C}$ and any $p\in E$, the relation
    $\Theta(\lambda p)=\overline{\lambda}\Theta(p)$
    holds.
  3. (3)
    The map $\Theta$ satisfies the involution property, i.e.,
    $\Theta^{2}=1_{E},$
    where $1_{E}$ denotes the identity map on $E$.
The pair $(E,\Theta)$ satisfying these conditions is called a real vector bundle.
Definition 5.3.
Let us consider two real vector bundles $(E,\Theta)$ and $(E^{\prime},\Theta^{\prime})$ over the same involution space $(X,\tau)$. We define an isomorphism between real vector bundles as a vector bundle isomorphism $f\colon E\to E^{\prime}$ that commutes with the involutions, meaning that the following equation holds:
$f\circ\Theta=\Theta^{\prime}\circ f.$
The isomorphism class of real vector bundles is the set of equivalence classes determined by real isomorphisms, which we denote by $\text{Vect}_{\mathcal{R}}^{m}(X,\tau)$, where $m$ represents the dimension of the fibers. We define $\mathrm{Vect}_{\mathcal{R}}(X,\tau)$ as the union of $\mathrm{Vect}_{\mathcal{R}}^{m}(X,\tau)$ over all non-negative integers $m$:
$\mathrm{Vect}_{\mathcal{R}}(X,\tau)=\bigcup_{m\geq 0}\mathrm{Vect}_{\mathcal{R}}^{m}(X,\tau).$
We consider a map from the isomorphism classes of real vector bundles $\text{Vect}_{\mathcal{R}}^{m}(X,\tau)$ to the isomorphism classes of complex vector bundles $\text{Vect}_{\mathbb{C}}^{m}(X)$. We define the operation of forgetting the real structure and denote this map as
$J:\text{Vect}_{\mathcal{R}}^{m}(X,\tau)\to\text{Vect}_{\mathbb{C}}^{m}(X).$
This map sends a real vector bundle $(E,\Theta)$ to its underlying complex vector bundle $E$, disregarding $\Theta$. The map $J$ is well-defined, as a real isomorphism includes a complex vector bundle isomorphism, and the forgetting operation consistently determines the isomorphism class.
Definition 5.4.
We consider two involution spaces $(X_{1},\tau_{1})$ and $(X_{2},\tau_{2})$.
  1. (1)
    A map $\varphi:X_{1}\to X_{2}$ is $\mathbb{Z}_{2}$-equivariant if the condition $\varphi\circ\tau_{1}=\tau_{2}\circ\varphi$ holds.
  2. (2)
    A $\mathbb{Z}_{2}$-equivariant homotopy between $\mathbb{Z}_{2}$-equivariant maps $\varphi_{0},\varphi_{1}:X_{1}\to X_{2}$ is defined as a continuous map $F:[0,1]\times X_{1}\to X_{2}$ such that for each $t\in[0,1]$, the map $\varphi_{t}(x)=F(t,x)$ is $\mathbb{Z}_{2}$-equivariant. We define the $\mathbb{Z}_{2}$-equivariant homotopy set $[X_{1},X_{2}]_{\text{eq}}$ as the set of equivalence classes of $\mathbb{Z}_{2}$-equivariant maps identified by $\mathbb{Z}_{2}$-equivariant homotopy.
Definition 5.5.
Let us consider the Grassmannian $G_{m}(\mathbb{C}^{\infty})$. We define the universal bundle $\mathcal{U}_{m}^{\infty}\to G_{m}(\mathbb{C}^{\infty})$ as follows. The total space $\mathcal{U}_{m}^{\infty}$ consists of pairs $(\Sigma,v)$, where $\Sigma\in G_{m}(\mathbb{C}^{\infty})$ is an $m$-dimensional subspace and $v\in\Sigma$ is a vector in that subspace. The projection map is given by
$\pi\colon\mathcal{U}_{m}^{\infty}\to G_{m}(\mathbb{C}^{\infty}),\quad(\Sigma,v)\mapsto\Sigma.$
We equip the universal bundle with a real structure. Specifically, we define an anti-linear homeomorphism $\Xi\colon\mathcal{U}_{m}^{\infty}\to\mathcal{U}_{m}^{\infty}$ by
$\Xi\colon(\Sigma,v)\mapsto(\varrho(\Sigma),\overline{v}),$
where $\varrho$ is the involution on $G_{m}(\mathbb{C}^{\infty})$. This map satisfies the following conditions:
  • The projection is compatible with the involution, i.e., $\pi\circ\Xi=\varrho\circ\pi$,
  • The map $\Xi$ is an involution, i.e., $\Xi^{2}=1_{\mathcal{U}_{m}^{\infty}}$,
where $1_{\mathcal{U}_{m}^{\infty}}$ denotes the identity map on $\mathcal{U}_{m}^{\infty}$. Thus, the pair $(\mathcal{U}_{m}^{\infty},\Xi)$ forms a real vector bundle.
The space $BU$ is equipped with the induced involution $\varrho$. We consider the universal bundle $\mathcal{U}^{\infty}\to BU$, which is endowed with a real structure $\Xi$ extending the structure on each $G_{m}(\mathbb{C}^{\infty})$. This real structure ensures that $(\mathcal{U}^{\infty},\Xi)$ forms a real vector bundle.
Theorem 5.6.
Let $(X,\tau)$ be an involution space, where $X$ is a compact Hausdorff space. For each $p\geq 0$, there exists a bijection
$[X,BU(p)]_{\text{eq}}\to\text{Vect}_{\mathcal{R}}^{p}(X,\tau),$
given by the pullback map
$f\mapsto f^{*}(\mathcal{U}_{p}^{\infty}).$
Moreover, there exists a bijection
$[X,BU]_{\text{eq}}\to\text{Vect}_{\mathcal{R}}(X,\tau),$
given by the pullback map
$f\mapsto f^{*}(\mathcal{U}^{\infty}).$
Proof.
See [26]. ∎
5.2. Proof of the bulk-edge correspondence
Here we give a proof of the bulk-edge correspondence for four-dimensional class AI topological insulators. The idea of the proof originates from [14].
Let $\hat{H}\colon T^{4}\to Herm(\mathbb{C}^{r})^{*}$ be a continuous map from the 4-dimensional torus $T^{4}=\mathbb{R}^{4}/2\pi\mathbb{Z}^{4}$ to the space $Herm(\mathbb{C}^{r})^{*}$ of invertible $r$ by $r$ Hermitian matrices. We assume that $\hat{H}(k)$ has both positive and negative eigenvalues at each $k\in T^{4}$. Associated to $\hat{H}$ is the Bloch vector bundle $E_{\hat{H}}\to T^{4}$ whose fiber at $k\in T^{4}$ consists of negative eigenvectors of $\hat{H}(k)$
$E_{\hat{H}}=\bigcup_{k\in T^{4}}\bigoplus_{\lambda<0}\ker(\hat{H}(k)-\lambda)\subset T^{4}\times\mathbb{C}^{r}.$
By evaluating the fundamental class $[T^{4}]\in H_{4}(T^{4};\mathbb{Z})$ of $T^{4}$ by the second Chern class $c_{2}(E_{\hat{H}})\in H^{4}(T^{4};\mathbb{Z})$ of Bloch bundle, we get the second Chern number of $E_{\hat{H}}$, which we denote by
$c_{2}(\hat{H})=\langle c_{2}(E_{\hat{H}}),[T^{4}]\rangle\in\mathbb{Z}.$
We define $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and identify $T^{4}=S^{1}\times S^{1}\times S^{1}\times S^{1}$. For $(k_{1},k_{2},k_{3})\in T^{3}=S^{1}\times S^{1}\times S^{1}$, we define $\hat{H}^{\sharp}(k_{1},k_{2},k_{3}):l^{2}(\mathbb{N},\mathbb{C}^{r})\to l^{2}(\mathbb{N},\mathbb{C}^{r})$ to be the compression of the multiplication operator with $\hat{H}(k_{1},k_{2},k_{3},\cdot):T^{3}\to Herm(\mathbb{C}^{r})^{*}$, namely, the composition of
$l^{2}(\mathbb{N},\mathbb{C}^{r})\overset{\hat{P}^{*}}{\longrightarrow}l^{2}(\mathbb{Z},\mathbb{C}^{r})\overset{\hat{H}(k_{1},k_{2},k_{3},\cdot)}{\longrightarrow}l^{2}(\mathbb{Z},\mathbb{C}^{r})\overset{\hat{P}}{\longrightarrow}l^{2}(\mathbb{N},\mathbb{C}^{r}),$
where $\hat{P}$ is the orthogonal projection. Notice that $\hat{H}^{\sharp}(k_{1},k_{2},k_{3})$ is essentially the Toeplitz operator [12] associated to $\hat{H}(k_{1},k_{2},k_{3},\cdot).$ Then $\hat{H}^{\sharp}(k_{1},k_{2},k_{3})$ is a self-adjoint bounded Fredholm operator. By the assumption that $\hat{H}(k_{1},k_{2},k_{3},k_{4})$ has both positive and negative eigenvalues, we obtain a continuous map $\hat{H}^{\sharp}:S^{1}\to Fred_{sa}^{1}$.
We define the involution $\tau$ on the torus $T^{4}$ as follows:
$\tau(k)=-k.$
We assume that $\hat{H}$ has time-reversal symmetry corresponding to class AI. That is, we assume that $K\hat{H}_{x}K=\hat{H}_{\tau(x)}$ for any $x\in T^{4}$. Here, $K$ is the complex conjugation acting on $\mathbb{C}^{r}$. That is, for $(a_{1},a_{2},\dots,a_{r})\in\mathbb{C}^{r}$,
$K(a_{1},a_{2},\dots,a_{r})=(\bar{a_{1}},\bar{a_{2}},\dots,\bar{a_{r}}).$
In this case, the Bloch bundle $E_{\hat{H}}$ becomes a real vector bundle in the sense of Atiyah [4]. The Bloch bundle $E_{\hat{H}}$ is a real vector bundle because $K$ induces an antilinear map between the fibers over $x\in X$ and $\tau(x)\in X$.
Theorem 5.7.
If $\hat{H}:T^{4}\to Herm(\mathbb{C}^{r})^{*}$ is a continuous map with time-reversal symmetry of class AI and $\hat{H}(k)$ has positive and negative eigenvalues at each $k\in T^{4}$, then $c_{2}(\hat{H})=\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}^{\sharp}])$.
The proof of Theorem 5.7 will be given below.
Tentatively, we denote by $Herm_{p}(\mathbb{C}^{r})$ the space of invertible Hermitian matrices which have $p$ negative eigenvalues and $r-p$ positive eigenvalues.
Lemma 5.8.
The following hold:
  1. (1)
    The spaces $\text{Herm}_{p}(\mathbb{C}^{r})$ and $\text{Gr}_{p}(\mathbb{C}^{r})$ are $\mathbb{Z}_{2}$-equivariantly homotopy equivalent under the actions $\tau(H)=\overline{H}$ and $\varrho(V)=\overline{V}$.
  2. (2)
    For a $\mathbb{Z}_{2}$-equivariant map $\hat{H}:T^{4}\to\text{Herm}_{p}(\mathbb{C}^{r})$, the pullback of the universal bundle $\mathcal{U}_{p}^{\infty}$ via this homotopy equivalence corresponds to the Bloch bundle $E_{\hat{H}}$, i.e., $h\circ\Phi_{*}([\hat{H}])=[E_{\hat{H}}]$, where $\Phi_{*}:[T^{4},\text{Herm}_{p}(\mathbb{C}^{r})]_{\mathbb{Z}_{2}}\to[T^{4},\text{Gr}_{p}(\mathbb{C}^{r})]_{\mathbb{Z}_{2}}$ is the map induced by the homotopy equivalence and $h:[T^{4},\text{Gr}_{p}(\mathbb{C}^{r})]_{\mathbb{Z}_{2}}\to\text{Vect}_{\mathfrak{R}}^{p}(T^{4})$ is defined by $h([g])=[g^{*}\mathcal{U}_{p}^{\infty}])$.
Proof.
We define $\Phi:\text{Herm}_{p}(\mathbb{C}^{r})\to\text{Gr}_{p}(\mathbb{C}^{r})$ by $\Phi(H)=\ker(1+\frac{H}{\|H\|})$, where the norm $\|H\|$ is the maximum absolute value of the eigenvalues of $H$, given by $\|H\|=\max(|\lambda_{\text{max}}(H)|,|\lambda_{\text{min}}(H)|)$, with $\lambda_{\text{max}}(H)$ being the largest positive eigenvalue and $\lambda_{\text{min}}(H)$ the smallest negative eigenvalue of $H$. We also define $\Psi:\text{Gr}_{p}(\mathbb{C}^{r})\to\text{Herm}_{p}(\mathbb{C}^{r})$ by $\Psi(V)=-P_{V}+P_{V^{\perp}}$, where $P_{V}$ is the orthogonal projection onto the $p$-dimensional subspace $V\subset\mathbb{C}^{r}$ and $P_{V^{\perp}}$ is the orthogonal projection onto its orthogonal complement $V^{\perp}$.
We first verify the $\mathbb{Z}_{2}$-equivariance of these maps. For $\Phi$,
$\Phi(\tau(H))$ $=\ker\left(1+\frac{\overline{H}}{\|\overline{H}\|}\right)=\overline{\ker\left(1+\frac{H}{\|H\|}\right)}=\varrho(\Phi(H)),$
and for $\Psi$,
$\Psi(\varrho(V))$ $=-P_{\overline{V}}+P_{\overline{V}^{\perp}}=\overline{-P_{V}+P_{V^{\perp}}}=\tau(\Psi(V)).$
To establish that $\Psi$ is a homotopy inverse of $\Phi$, that is, $\Phi\circ\Psi\simeq\mathrm{id}$ and $\Psi\circ\Phi\simeq\mathrm{id}$, we compute $\Phi\circ\Psi(V)$ as follows:
$\Phi\circ\Psi(V)=\Phi(-P_{V}+P_{V^{\perp}})=\ker(1+(-P_{V}+P_{V^{\perp}})).$
Since $-P_{V}+P_{V^{\perp}}$ has eigenvalue $-1$ on $V$ and eigenvalue $1$ on $V^{\perp}$, the operator $1+(-P_{V}+P_{V^{\perp}})$ has kernel equal to $V$. Therefore, we obtain $\Phi\circ\Psi(V)=V$. Thus, $\Phi\circ\Psi=1_{\text{Gr}_{p}(\mathbb{C}^{r})}$. For $\Psi\circ\Phi$, we define a homotopy
$H_{t}=(1-t)H+t(-P_{V_{H}}+P_{V_{H}^{\perp}}),\quad t\in[0,1],$
where $V_{H}=\Phi(H)$. The matrix $H_{t}$ remains Hermitian with $p$ negative eigenvalues throughout $t\in[0,1]$, and it satisfies $\tau(H_{t})=\overline{H_{t}}$, ensuring that $H_{t}$ lies in $\text{Herm}_{p}(\mathbb{C}^{r})$. This homotopy connects $H$ to $\Psi(\Phi(H))$, proving that $\Psi\circ\Phi\simeq 1_{\text{Herm}_{p}(\mathbb{C}^{r})}$. Hence, $\Psi$ is a $\mathbb{Z}_{2}$-equivariant homotopy inverse of $\Phi$.
Next, we consider a $\mathbb{Z}_{2}$-equivariant map $\hat{H}:T^{4}\to\text{Herm}_{p}(\mathbb{C}^{r})$ satisfying $\hat{H}(\tau(k))=\overline{\hat{H}(k)}$. The induced map $\Phi_{*}([\hat{H}])=[\Phi\circ\hat{H}]$ maps to $[T^{4},\text{Gr}_{p}(\mathbb{C}^{r})]_{\mathbb{Z}_{2}}$, and we compute $h([\Phi\circ\hat{H}])=[(\Phi\circ\hat{H})^{*}\mathcal{U}^{\infty}]$. Since $\mathcal{U}^{\infty}_{p}$ over $\text{Gr}_{p}(\mathbb{C}^{r})$ corresponds to the tautological bundle $\{(V,v)\mid v\in V\}$ under the equivalence $\text{Gr}_{p}(\mathbb{C}^{r})\simeq BU(p)$, the pullback is
$(\Phi\circ\hat{H})^{*}\mathcal{U}^{\infty}_{p}=\{(k,v)\mid v\in V_{\hat{H}(k)}\},$
where $V_{\hat{H}(k)}=\ker(1+\frac{\hat{H}(k)}{\|\hat{H}(k)\|})$ is the negative eigenspace of $\hat{H}(k)$. The Bloch bundle $E_{\hat{H}}$ is defined as
$E_{\hat{H}}=\{(k,v)\mid\hat{H}(k)v=\lambda(k)v,\lambda(k)<0\},$
which coincides with $V_{\hat{H}(k)}$ by the construction of $\Phi$. Therefore, we have $[(\Phi\circ\hat{H})^{*}\mathcal{U}^{\infty}_{p}]=[E_{\hat{H}}]$, and it follows that $h\circ\Phi_{*}([\hat{H}])=[E_{\hat{H}}]$. ∎
Lemma 5.9.
For any integer $n\in\mathbb{Z}$, we define $\hat{H}_{n}:T^{4}\to\mathbb{C}^{4}$ by
$\hat{H}_{n}(k_{1},k_{2},k_{3},k_{4})=\begin{pmatrix}a(k)&\bar{c}(k)-e^{ik_{4}}&0&\bar{b}(k)\\ c(k)-e^{-ik_{4}}&-a(k)&\bar{b}(k)&0\\ 0&b(k)&a(k)&-c(k)+e^{-ik_{4}}\\ b(k)&0&-\bar{c}(k)+e^{ik_{4}}&-a(k)\end{pmatrix}$
where
$a(k)=-\cos(k_{2}+k_{3}),$
$b(k)=-e^{-\sqrt{-1}nk_{1}}-e^{-\sqrt{-1}k_{2}},$
$c(k)=-1-e^{-\sqrt{-1}k_{3}}.$
Then, $c_{2}(\hat{H}_{n})=-2n$.
Proof.
In the case that $n=1$, according to the results of previous research [31], $c_{2}(\hat{H})=-2$. The map $H_{n}$ factors as $\hat{H}_{n}=\hat{H}_{1}\circ f_{n}$, where $f_{n}:T^{4}\to T^{4}$ is given by $f_{n}(k_{1},k_{2},k_{3},k_{4})=(nk_{1},k_{2},k_{3},k_{4})$. Since $f_{n}$ carries the fundamental class $[T^{4}]\in H_{4}(T^{4};\mathbb{Z})\cong\mathbb{Z}$ of the torus to $n[T^{4}]$, it follows that $c_{2}(\hat{H}_{n})=-2n$. ∎
Lemma 5.10.
For each $m\geq 2$ there exsits a bijection
$\text{Vect}_{\mathcal{R}}^{m}(T^{4},\tau)\cong 2\mathbb{Z}$
provided by the map $c_{2}\circ J$.
Proof.
Refer to [11]. ∎
Lemma 5.11.
Let $2\leq p<r$ and let $k$ be a sufficiently large natural number. Then, there exists an isomorphism
$[T^{4},\mathrm{Herm}_{p}(\mathbb{C}^{r+k})]_{eq}\cong 2\mathbb{Z},$
which is given by the map
$\hat{H}\mapsto c_{2}(E_{\hat{H}}).$
Proof.
We define
$\hat{F}_{n}(k)$ $=\hat{H}_{n}(k)\oplus-1_{\mathbb{C}^{p-2}}\oplus 1_{\mathbb{C}^{(r+k)-p-2}}$
by using $\hat{H}_{n}$ as specified in Lemma 5.9, where $1$ denotes the identity map. Since Lemma 5.9 shows that
$c_{2}(\hat{F}_{n})=-2n,$
the claim follows from Lemma 5.8, Lemma 5.10, and Theorem 5.6. ∎
Lemma 5.12.
For $n\in\mathbb{Z}$, we have $\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}_{n}^{\sharp}]=-2n$
Proof.
By solving the equations $a(k)=0$, $b(k)=0$, and $|c(k)|\leq 0$ derived from Proposition 6.8, we can identify the Fermi set as follows:
$F=\bigcup_{l=1}^{n}\left\{\left(\frac{2\pi}{3n}+\frac{2(l-1)\pi}{n},\frac{4\pi}{3},\frac{7\pi}{6}\right),\left(\frac{4\pi}{3n}+\frac{2(l-1)\pi}{n},\frac{2\pi}{3},\frac{5\pi}{6}\right)\right\}.$
Consequently, from Theorem 4.19, we obtain
$\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}_{n}^{\sharp}])=-2n.$
We are now in the position to prove Theorem 5.7:
Proof.
Because of Lemma 5.11, if a map $\hat{H}:T^{4}\to Herm(\mathbb{C}^{r})^{*}$ satisfying the hypothesis of Theorem 5.7 has the Chern number $c_{2}(\hat{H})=-2n$, then for a sufficiently large natural number $k$, $\hat{H}\oplus 1_{\mathbb{C}^{k}}$ is homotopic to $\hat{F}_{n}$. Here, $\hat{F}_{n}$ is the map defined in the proof of Lemma 5.11. Then, the associated families of self-adjoint Fredholm operators $(\hat{H}\oplus 1_{\mathbb{C}^{k}})^{\sharp}:T^{3}\to Fred_{sa}^{1}$, $\hat{H}^{\sharp}:T^{3}\to Fred_{sa}^{1}$, and $\hat{F}^{\sharp}_{n}:T^{3}\to Fred_{sa}^{1}$ are also homotopic. Since the Chern character is a homotopy invariant, it follows that
$\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([(\hat{H}\oplus 1_{\mathbb{C}^{k}})^{\sharp}])=\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}(\hat{F}^{\sharp}_{n}).$
By Lemma 5.12, we see
$\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}(\hat{F}^{\sharp}_{n})=\int_{T^{3}}\operatorname{ch}_{\frac{3}{2}}([\hat{H}_{n}^{\sharp}])=-2n.$
Hence $c_{2}(\hat{H})=\int_{T^{3}}ch_{\frac{3}{2}}([\hat{H}_{n}^{\sharp}])$ is established. ∎
6. Spectral analysis of the local model
Let $\hat{H}_{(a,b,c)}\colon S^{1}\to Herm(\mathbb{C}^{4})^{*}$ be the continuous map given by
$\hat{H}_{(a,b,c)}(k)=\hat{H}_{loc}(a,b,c,k)=\begin{pmatrix}a&\bar{c}-e^{ik}&0&\bar{b}\\ c-e^{-ik}&-a&\bar{b}&0\\ 0&b&a&-c+e^{-ik}\\ b&0&-\bar{c}+e^{ik}&-a\end{pmatrix}$
where $a\in\mathbb{R}$ and $b,c\in\mathbb{C}.$ Let $H_{(a,b,c)}\colon l^{2}(\mathbb{Z},\mathbb{C}^{4})\to l^{2}(\mathbb{Z},\mathbb{C}^{4})$ be the Fourier transformation of the multiplication operator with $\hat{H}$. Concretely, $H_{(a,b,c)}$ acts on $\psi=(\psi(n))_{n\in\mathbb{Z}}\in l^{2}(\mathbb{Z},\mathbb{C}^{4})$ by
$(H_{(a,b,c)}\psi)(n)=A\psi(n-1)+V\psi(n)+A^{*}\psi(n+1),$
in which the matrices $A$ and $V$ are
$A$ $=\begin{pmatrix}0&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&1&0\end{pmatrix},$ $V$ $=\begin{pmatrix}a&\bar{c}&0&\bar{b}\\ c&-a&\bar{b}&0\\ 0&b&a&-c\\ b&0&-\bar{c}&-a\end{pmatrix}.$
Define $H_{(a,b,c)}^{\#}\colon l^{2}(\mathbb{N},\mathbb{C}^{4})\to l^{2}(\mathbb{N},\mathbb{C}^{4})$ by $PH_{(a,b,c)}P^{*}$, where $P\colon l^{2}(\mathbb{Z},\mathbb{C}^{4})\to l^{2}(\mathbb{N},\mathbb{C}^{4})$ is the orthogonal projection. Hereafter, $H^{\#}_{(a,b,c)}$ will be simply written as $H^{\#}$.
Given $E\in\mathbb{R}$, we are to consider the solutions to the equation $(H^{\#}-E)\psi=0$, which is equivalent to the equations for $\psi(1),\psi(2),\psi(3),\ldots\in\mathbb{C}^{4},$
$0=(V-E)\psi(1)+A^{*}\psi(2),$
$0=A\psi(1)+(V-E)\psi(2)+A^{*}\psi(3),$
$0=A\psi(2)+(V-E)\psi(3)+A^{*}\psi(4),$
$\vdots$
$0=A\psi(n-1)+(V-E)\psi(n)+A^{*}\psi(n+1).$
Also, for any $n\in\mathbb{N}$, we define $e_{1},e_{4}\in l^{2}(\mathbb{N},\mathbb{C}^{4})$ by $e_{1}(n)={}^{t}\!(c^{n-1},0,0,0)\in\mathbb{C}^{4},e_{4}(n)={}^{t}\!(0,0,0,c^{n-1})\in\mathbb{C}^{4}$ Here, we assume that the complex parameter $c$ appearing in the definition of $H_{(a,b,c)}$ satisfies $|c|<1$, and we adopt the convention that $10^{0}=$ 1. In particular, whenever we consider the vectors $e_{1}$ and $e_{4}$, we do so under the assumption that $|c|<1$, where $c\in\mathbb{C}$ is the same parameter as in $H_{(a,b,c)}$. Under this convention, for $c=0$, we have
$e_{1}(1)={}^{t}\!(1,0,0,0),\quad e_{1}(n)={}^{t}\!(0,0,0,0)\ (n\geq 2),$
and
$e_{4}(1)={}^{t}\!(0,0,0,1),\quad e_{4}(n)={}^{t}\!(0,0,0,0)\ (n\geq 2).$
Let $V$ be a complex vector space. In the following, for any $x_{1},x_{2},\ldots,x_{n}\in V$, we write the subspace generated by $x_{1},x_{2},\ldots,x_{n}\in V$ as $\langle x_{1},x_{2},...,x_{n}\rangle\subset V.$
Lemma 6.1.
Suppose that $b=c=0$.
  1. (1)
    In the case that $E=\pm\sqrt{a^{2}+1}$, the equation $(H^{\#}-E)\psi=0$ admits $l^{2}$-solutions. The solutions form a vector space of infinite dimension.
  2. (2)
    In the case that $E=\pm a$, the equation $(H^{\#}-E)\psi=0$ admits $l^{2}$-solutions. In this case, the solution space is given as follows:
    $\ker(H^{\#}-E)$ =$=\begin{cases}\langle e_{1}\rangle&E=a\neq 0.\\ \langle e_{4}\rangle&E=-a\neq 0.\\ \langle e_{1},e_{4}\rangle&E=a=0.\end{cases},$ $\dim(\ker(H^{\#}-E))$ =$=\begin{cases}1&a\neq 0.\\ 2&a=0.\end{cases}$
  3. (3)
    In the other cases, the equation $(H^{\#}-E)\psi=0$ admits no non-trivial $l^{2}$-solutions.
Proof.
We express $\psi(n)={}^{t}\!(\alpha(n),\beta(n),\gamma(n),\delta(n)).$ The equation $0=(V-E)\psi(1)+A^{*}\psi(2)$ is equivalent to
$(a-E)\alpha(1)$ $=0,$ $-(a+E)\beta(1)-\alpha(2)$ $=0,$
$(a-E)\gamma(1)+\delta(2)$ $=0,$ $-(a+E)\delta(1)$ $=0.$
For $n\geq 1$, the components of $0=A\psi(n)+(V-E)\psi(n+1)+A^{*}\psi(n+2)$ are
$-\beta(n)+(a-E)\alpha(n+1)$ $=0,$ $-(a+E)\beta(n+1)-\alpha(n+2)$ $=0,$
$(a-E)\gamma(n+1)+\delta(n+2)$ $=0,$ $\gamma(n)-(a+E)\delta(n+1)$ $=0.$
  1. (1)
    We consider the case where the condition $a^{2}-E^{2}\neq 0$ holds. Under this assumption, the following relations are satisfied for all $n\geq 1$:
    $\frac{1}{a-E}\beta(n)$ $=-(a+E)\beta(n),$ $\frac{-1}{a-E}\delta(n+1)$ $=(a+E)\delta(n+1).$
    If $\beta(n)=0$ and $\delta(n+1)=0$ for all $n\in\mathbb{N}$, then the solution to the equation $(H^{\#}-E)\psi=0$ is the trivial solution. However, if there exists some $n_{0}\in\mathbb{N}$ such that $\beta(n_{0})\neq 0$ or $\delta(n_{0}+1)\neq 0$, then the eigenvalue $E$ satisfies:
    $E=\pm\sqrt{a^{2}+1}.$
    Under this derived condition on $E$, the following equations hold for all $n\geq 2$:
    $\alpha(n+1)$ $=-(a+E)\beta(n),$ $\delta(n+1)$ $=-(a+E)\gamma(n),$ $\alpha(1)$ $=0,$ $\delta(1)$ $=0,$
    and for any sequences $\beta=(\beta(n))_{n\in\mathbb{N}}$ and $\gamma=(\gamma(n))_{n\in\mathbb{N}}$ in $l^{2}(\mathbb{N},\mathbb{C}^{4})$, the norm of the solution $\psi$ is finite:
    $\lVert\psi\rVert^{2}=[1+(a+E)^{2}]\left(\sum_{n\geq 1}|\beta(n)|^{2}+\sum_{n\geq 1}|\gamma(n)|^{2}\right)<\infty.$
    Consequently, the dimension of the kernel of $H^{\#}-E$ is infinite, i.e., $\dim(\ker(H^{\#}-E))=\infty$.
  2. (2)
    We consider the case where $E=a$. Under this assumption, the components of the equation $(H^{\#}-E)\psi=0$ for all $n\geq 2$ are given by:
    $\beta(n-1)$ $=0,$ $-2a\beta(n)-\alpha(n+1)$ $=0,$
    $\delta(n+1)$ $=0,$ $\gamma(n-1)-2a\delta(n)$ $=0.$
    Additionally, for $n=1$, the components of the equation $(H^{\#}-E)\psi=0$ are:
    $\beta(1)$ $=0,$ $-2a\beta(1)-\alpha(2)$ $=0,$ $\delta(2)$ $=0.$
    From these equations, we deduce that $\alpha(n)=0$ holds for all $n\geq 2$, and $\gamma(n)=0$ holds for all $n\geq 1$. Furthermore, the equation $a\delta(1)=0$ must be satisfied for $n=1$. Consequently, the value of $\delta(1)$ depends on $a$ as follows: if $a=0$, then $\delta(1)$ can be any complex number, whereas if $a\neq 0$, then $\delta(1)=0$.
    Thus, the kernel of $H^{\#}-E$ is determined as:
    $\ker(H^{\#}-E)=\begin{cases}\langle e_{1}\rangle&\text{if }a\neq 0,\\ \langle e_{1},e_{4}\rangle&\text{if }a=0.\end{cases}$
  3. (3)
    We consider the case where $E=-a\neq 0$. Under this assumption, the components of the equation $H^{\#}-E=0$ for all $n\geq 2$ are given by:
    $-\beta(n-1)+2a\alpha(n)$ $=0,$ $\alpha(n+1)$ $=0,$ $2a\gamma(n)+\delta(n+1)$ $=0,$ $\gamma(n-1)$ $=0.$
    For $n=1$, the components of the equation $H^{\#}-E=0$ are:
    $2a\alpha(1)$ $=0,$ $\alpha(2)$ $=0,$
    $2a\gamma(1)+\delta(2)$ $=0,$ $\gamma(1)$ $=0.$
    From these equations, we deduce that $\alpha(n)=0$ and $\delta(n)=0$ hold for all $n\geq 2$. Additionally, we find that $\beta(n)=0$ and $\gamma(n)=0$ hold for all $n\geq 1$. Consequently, $\delta(1)$ can be any complex number, and since the condition $a\neq 0$ holds, $\alpha(1)=0$ follows from the equation $02a\alpha(1)=$ 0.
    Thus, the kernel of $H^{\#}-E$ is given by:
    $\ker(H^{\#}-E)=\langle e_{4}\rangle.$
Lemma 6.2.
Suppose that $b\neq 0$ and $c=0$.
  1. (1)
    In the case that $E=\pm\sqrt{a^{2}+|b|^{2}+1}$, the equation $(H^{\#}-E)\psi=0$ admits $l^{2}$-solutions. The solutions form a vector space of infinite dimension.
  2. (2)
    In the case that $E=\pm\sqrt{a^{2}+|b|^{2}}$, the equation $(H^{\#}-E)\psi=0$ admits $l^{2}$-solutions. In this case, one of the eigenvectors is $\frac{a+E}{b}e_{1}+e_{4}$, and the following holds:
    $\dim(\ker(H^{\#}-E))=1,$
    $\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}}))=\langle e_{1},e_{4}\rangle.$
  3. (3)
    In the other cases, the equation $(H^{\#}-E)\psi=0$ admits no non-trivial $l^{2}$-solutions.
Proof.
We examine the equation $(H^{\#}-E)\psi=0$. For all $n\geq 1$, the components of this equation are given by:
$-\beta(n-1)+(a-E)\alpha(n)+\bar{b}\delta(n)$ $=0,$
$-(a+E)\beta(n)+\bar{b}\gamma(n)-\alpha(n+1)$ $=0,$
$b\beta(n)+(a-E)\gamma(n)+\delta(n+1)$ $=0,$
$\gamma(n-1)+b\alpha(n)-(a+E)\delta(n)$ $=0.$
We formally define $\alpha(0)=\beta(0)=\delta(0)=\gamma(0)=0$. From the components of the equation $(H^{\#}-E)\psi=0$, we derive the following expressions:
$\alpha(n+1)$ $=-(a+E)\beta(n)+\bar{b}\gamma(n),$
$\delta(n+1)$ $=-b\beta(n)-(a-E)\gamma(n).$
By substituting these into the original equations and simplifying, we obtain the following relations for all $n\geq 1$:
$(a^{2}+|b|^{2}+1-E^{2})\beta(n)$ $=0,$
$(a^{2}+|b|^{2}+1-E^{2})\gamma(n)$ $=0.$
  1. (1)
    Suppose that the condition $a^{2}+|b|^{2}+1-E^{2}\neq 0$ holds. Under this assumption, $\beta(n)=0$ and $\gamma(n)=0$ must be true for all $n\in\mathbb{N}$. From the equations above, we deduce that $\alpha(n)=0$ and $\delta(n)=0$ hold for all $n\geq 2$. Additionally, for $n=1$, the remaining components yield:
    $(a-E)\alpha(1)+\bar{b}\delta(1)$ $=0,$
    $b\alpha(1)-(a+E)\delta(1)$ $=0.$
    We then analyze the solutions to these equations. If the condition $a^{2}-E^{2}+|b|^{2}\neq 0$ is satisfied, then only the trivial solution $\alpha(1)=0$ and $\delta(1)=0$ exists. However, if $a^{2}-E^{2}+|b|^{2}=0$, the equations imply:
    $\alpha(1)-\frac{a+E}{b}\delta(1)$ $=0.$
    Thus, when the eigenvalue $E$ satisfies $E=\pm\sqrt{a^{2}+|b|^{2}}$, the kernel of $H^{\#}-E$ is given by:
    $\ker(H^{\#}-E)=\left\langle\frac{a+E}{b}e_{1}+e_{4}\right\rangle.$
  2. (2)
    Suppose that the condition $a^{2}+|b|^{2}+1-E^{2}=0$ holds. Under this assumption, the equations $(a^{2}+|b|^{2}+1-E^{2})\beta(n)=0$ and $(a^{2}+|b|^{2}+1-E^{2})\gamma(n)=0$ are trivially satisfied for any sequences $(\beta(n))_{n\in\mathbb{N}}$ and $(\gamma(n))_{n\in\mathbb{N}}$ in $l^{2}(\mathbb{N},\mathbb{C})$. In this case, the sequences $(\alpha(n))_{n\in\mathbb{N}}$ and $(\delta(n))_{n\in\mathbb{N}}$, defined by the derived expressions, are also elements of $l^{2}(\mathbb{N},\mathbb{C})$. Therefore, the solution $\psi$ belongs to $l^{2}(\mathbb{N},\mathbb{C}^{4})$, and the dimension of the kernel of $H^{\#}-E$ is infinite, i.e., $\dim(\ker(H^{\#}-E))=\infty$.
Additionally, we express the basis vectors $e_{1}$ and $e_{4}$ as follows:
$e_{1}$ $=\frac{b}{2\sqrt{a^{2}+|b|^{2}}}\left[\left(\frac{a+\sqrt{a^{2}+|b|^{2}}}{b}e_{1}+e_{4}\right)-\left(\frac{a-\sqrt{a^{2}+|b|^{2}}}{b}e_{1}+e_{4}\right)\right],$
$\in\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}})),$
$e_{4}$ $=\left(\frac{a+\sqrt{a^{2}+|b|^{2}}}{b}e_{1}+e_{4}\right)-\frac{a+\sqrt{a^{2}+|b|^{2}}}{b}e_{1},$
$\in\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}})).$
Therefore, the direct sum of the kernels satisfies:
$\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}}))=\langle e_{1},e_{4}\rangle.$
Lemma 6.3.
Suppose that $c\neq 0$. If we ignore the $l^{2}$-condition, then any solution $\psi=(\psi(n))_{n\in\mathbb{N}}$ to $(H^{\#}-E)\psi=0$ can be described as follows
$\psi(1)=z\begin{pmatrix}\bar{c}\\ -(a-E)\\ b\\ 0\end{pmatrix}+w\begin{pmatrix}0\\ -\bar{b}\\ -(a+E)\\ \bar{c}\end{pmatrix},\quad$ $\psi(n+1)=R\psi(n),\quad(n\geq 1)$
where $z$ and $w$ are complex numbers, and
$R=\begin{pmatrix}c&-(a+E)&\bar{b}&0\\ -\frac{c(a-E)}{\bar{c}}&\frac{a^{2}+|b|^{2}+1-E^{2}}{\bar{c}}&0&-\frac{\bar{b}c}{\bar{c}}\\ \frac{bc}{\bar{c}}&0&\frac{a^{2}+|b|^{2}+1-E^{2}}{\bar{c}}&-\frac{(a+E)c}{\bar{c}}\\ 0&-b&-(a-E)&c\end{pmatrix}.$
Proof.
We express the vector $\psi(n)$ as $\psi(n)={}^{t}(\alpha(n),\beta(n),\gamma(n),\delta(n))$. The equation $0=(V-E)\psi(1)+A^{*}\psi(2)$ is equivalent to the following system of equations:
$(a-E)\alpha(1)+\bar{c}\beta(1)+\bar{b}\delta(1)$ $=0,$
$c\alpha(1)-(a+E)\beta(1)+\bar{b}\gamma(1)-\alpha(2)$ $=0,$
$b\beta(1)+(a-E)\gamma(1)-c\delta(1)+\delta(2)$ $=0,$
$b\alpha(1)-\bar{c}\gamma(1)-(a+E)\delta(1)$ $=0.$
Under the assumption that $c\neq 0$ holds, we solve these equations and obtain:
$\beta(1)$ $=-\frac{a-E}{\bar{c}}\alpha(1)-\frac{\bar{b}}{\bar{c}}\delta(1),$
$\gamma(1)$ $=\frac{b}{\bar{c}}\alpha(1)-\frac{a+E}{\bar{c}}\delta(1).$
For all $n\geq 2$, the equation $0=A\psi(n-1)+(V-E)\psi(n)+A^{*}\psi(n+1)$ leads to the following system:
$-\beta(n-1)+(a-E)\alpha(n)+\bar{c}\beta(n)+\bar{b}\delta(n)$ $=0,$
$c\alpha(n)-(a+E)\beta(n)+\bar{b}\gamma(n)-\alpha(n+1)$ $=0,$
$b\beta(n)+(a-E)\gamma(n)-c\delta(n)+\delta(n+1)$ $=0,$
$\gamma(n-1)+b\alpha(n)-\bar{c}\gamma(n)-(a+E)\delta(n)$ $=0.$
From this system, we derive the following recursive relations for all $n\geq 2$:
$\alpha(n+1)$ $=c\alpha(n)-(a+E)\beta(n)+\bar{b}\gamma(n),$
$\delta(n+1)$ $=-b\beta(n)-(a-E)\gamma(n)+c\delta(n).$
By substituting these relations into the system, we obtain the following expressions for all $n\geq 2$:
$\beta(n)$ $=\frac{1}{\bar{c}}\big{[}-c(a-E)\alpha(n-1)+(a^{2}+|b|^{2}+1-E^{2})\beta(n-1)-\bar{b}c\delta(n-1)\big{]},$
$\gamma(n)$ $=\frac{1}{\bar{c}}\big{[}bc\alpha(n-1)+(a^{2}+|b|^{2}+1-E^{2})\gamma(n-1)-(a+E)c\delta(n-1)\big{]}.$
Thus, we complete the proof of the desired result. ∎
We compute the determinant of the matrix $R$ and find that it is given by:
$\det(R)=\left(\frac{c}{\bar{c}}\right)^{2}.$
We also consider the quadratic equation in $\lambda$, defined by $\bar{c}\lambda^{2}-(a^{2}+|b|^{2}+|c|^{2}+1-E^{2})\lambda+c=0$, which arises from the characteristic polynomial $\det(\lambda I-R)=0$. The discriminant of this quadratic equation is:
$\Delta$ $=(a^{2}+|b|^{2}+|c|^{2}+1-E^{2})^{2}-4|c|^{2}$
$=(a^{2}+|b|^{2}+(|c|-1)^{2}-E^{2})(a^{2}+|b|^{2}+(|c|+1)^{2}-E^{2}).$
The two solutions to the equation $\det(\lambda I-R)=0$ are:
$\lambda_{1},\lambda_{2}=\frac{(a^{2}+|b|^{2}+|c|^{2}+1-E^{2})\mp\sqrt{\Delta}}{2\bar{c}}.$
Additionally, we define the quantities $a_{1}$ and $a_{2}$ as follows: $a_{1}=\bar{c}\lambda_{1}$ and $a_{2}=\bar{c}\lambda_{2}$.
Lemma 6.4.
Suppose that $c\neq 0$. In this case, if $\Delta\leq 0$, then there is no $l^{2}$-solutions to $(H^{\#}-E)\psi=0$.
Proof.
Since the discriminant $\Delta$ satisfies $\Delta\leq 0$, we conclude that the quantities $a_{1}$ and $a_{2}$ are complex numbers, and their complex conjugates satisfy $\overline{a_{2}}=a_{1}$. Consequently, we derive the relation $\lambda_{2}=\frac{\bar{c}}{c}\lambda_{1}$. Additionally, from the relationship between the roots and the coefficients of the characteristic polynomial $\det(\lambda I-R)=0$, we obtain $\lambda_{1}\lambda_{2}=\frac{c}{\bar{c}}$. Therefore, we find that the norm satisfies $|\lambda_{1}|=|\lambda_{2}|=1$, and in particular, $\lambda_{1}\neq 0$ and $\lambda_{2}\neq 0$.
  1. (1)
    Suppose that the matrix $R$ is diagonalizable. Under this assumption, there exists an invertible matrix $T\in GL(4,\mathbb{C})$ such that we can express $T^{-1}RT=\text{diag}(\lambda_{1},\lambda_{1},\lambda_{2},\lambda_{2})$. We note that a solution $\psi$ to the equation $(H^{\#}-E)\psi=0$ in the space $\text{Map}(\mathbb{N},\mathbb{C}^{4})$ satisfies $\psi\in l^{2}(\mathbb{N},\mathbb{C}^{4})$ if and only if $T^{-1}\psi\in l^{2}(\mathbb{N},\mathbb{C}^{4})$.
    We express the the vector $T^{-1}\psi(n)$ as $T^{-1}\psi(n)={}^{t}(x(n),y(n),z(n),w(n))$. Then, we compute the $l^{2}$-norm of $T^{-1}\psi$ as follows:
    $\|T^{-1}\psi\|^{2}$ $=\sum_{n\in\mathbb{N}}\big{(}|x(n)|^{2}+|y(n)|^{2}+|z(n)|^{2}+|w(n)|^{2}\big{)}$
    $=\sum_{n\in\mathbb{N}}\big{(}|\lambda_{1}|^{2n}(|x(1)|^{2}+|y(1)|^{2})+|\lambda_{2}|^{2n}(|z(1)|^{2}+|w(1)|^{2})\big{)}$
    $=\sum_{n\in\mathbb{N}}\big{(}|x(1)|^{2}+|y(1)|^{2}+|z(1)|^{2}+|w(1)|^{2}\big{)},$
    where the last equality holds because $|\lambda_{1}|=|\lambda_{2}|=1$. Therefore, if $\psi\neq 0$, we conclude that $\|T^{-1}\psi\|^{2}=\infty$.
  2. (2)
    Suppose that the matrix $R$ is not diagonalizable. Under this assumption, there exists an invertible matrix $T\in GL(4,\mathbb{C})$ such that we can express:
    $T^{-1}RT=\begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&0&0\\ 0&0&\lambda_{2}&1\\ 0&0&0&\lambda_{2}\end{pmatrix}.$
    As in the previous case, we analyze the $l^{2}$-norm of $T^{-1}\psi$. Unless $\psi=0$, the $l^{2}$-norm of $T^{-1}\psi$ diverges due to the polynomial growth induced by the Jordan blocks, and consequently, the norm $\|\psi\|$ also diverges.
Thus, we complete the proof. ∎
Lemma 6.5.
We suppose that the conditions $c\neq 0$, $b=0$, and $a=E$ hold. Under these assumptions, we have $\Delta>0$ if and only if $|c|\neq 1$. Furthermore, the equation $(H^{\#}-E)\psi=0$ has nontrivial $l^{2}$-solutions if and only if $|c|<1$. Additionally, when the condition $|c|<1$ is satisfied, we have the following properties of the kernel:
$\ker(H^{\#}-E)$ =$=\begin{cases}\langle e_{1}\rangle&a\neq 0,\\ \langle e_{1},e_{4}\rangle&a=0,\end{cases}$ $\dim(\ker(H^{\#}-E))$ =$=\begin{cases}1&a\neq 0,\\ 2&a=0.\end{cases}$
Proof.
We have $\Delta=(|c|^{2}-1)^{2}$, which is greater than zero if and only if the norm of $c$ is not equal to 1. Thus, we proceed under the assumption that $|c|\neq 1$ holds throughout the following discussion.
  1. (1)
    The condition $|c|<1$ is satisfied. In this case, we find that the eigenvalues are $\lambda_{1}=c$ and $\lambda_{2}=\frac{1}{\bar{c}}$. We introduce a matrix $T$ defined as follows:
    $T=\begin{pmatrix}0&1&0&\frac{2a\bar{c}}{|c|^{2}-1}\\ 0&0&0&1\\ -\frac{2ac}{|c|^{2}-1}&0&1&0\\ 1&0&0&0\end{pmatrix}.$
    Since the determinant of $T$ is $1$, we conclude that $T$ is an invertible matrix. The inverse is
    $T^{-1}=\begin{pmatrix}0&0&0&1\\ 1&-\frac{2a\bar{c}}{|c|^{2}-1}&0&0\\ 0&0&1&\frac{2ac}{|c|^{2}-1}\\ 0&1&0&0\end{pmatrix}.$
    Under this transformation, we verify that $T^{-1}RT=\text{diag}(c,c,\frac{1}{\bar{c}},\frac{1}{\bar{c}})$. We express the transformed solution $T^{-1}\psi(n)$ as the transpose $(x(n),y(n),z(n),w(n))$. For $n=1$, we obtain:
    $T^{-1}\psi(1)=z\begin{pmatrix}0\\ \bar{c}\\ 0\\ 0\end{pmatrix}+w\begin{pmatrix}\bar{c}\\ 0\\ \frac{2a}{|c|^{2}-1}\\ 0\end{pmatrix},$
    and we compute the $l^{2}$-norm of $T^{-1}\psi$ as:
    $\|T^{-1}\psi\|^{2}=\sum_{n\in\mathbb{N}}\left(|c|^{2n}\big{(}|x(n)|^{2}+|y(n)|^{2}\big{)}+\left|\frac{1}{\bar{c}}\right|^{2n}\big{(}|z(n)|^{2}+|w(n)|^{2}\big{)}\right).$
    Since $|\lambda_{1}||\lambda_{2}|=1$ and $|c|<1$, we conclude that $\|T^{-1}\psi\|^{2}<\infty$ holds if and only if either $z(1)=w(1)=0$ or both $x(1)=y(1)=0$ and $\left|\frac{1}{\bar{c}}\right|<1$.
    1. (a)
      We suppose that $z(1)=w(1)=0$ holds. Since $w(1)=w\frac{2a}{|c|^{2}-1}=0$, we deduce that either $w=0$ or $a=0$. Therefore, we determine that the kernel of $H^{\#}-E$ is:
      $\ker(H^{\#}-E)=\begin{cases}\langle e_{1}\rangle&\text{if }a\neq 0,\\ \langle e_{1},e_{4}\rangle&\text{if }a=0.\end{cases}$
    2. (b)
      We suppose that $\left|\frac{1}{\bar{c}}\right|<1$ and $x(1)=y(1)=0$ hold. Since $\left|\frac{1}{\bar{c}}\right|<1$ implies $|c|>1$, we recognize that this contradicts the assumption $|c|<1$ in this case.
  2. (2)
    We suppose that the conditions $\left|\frac{1}{\bar{c}}\right|<1$ and $x(1)=y(1)=0$ are satisfied, which implies $|c|>1$. In this case, we find that the eigenvalues are $\lambda_{1}=\frac{1}{\bar{c}}$ and $\lambda_{2}=c$. We introduce a matrix $T$ defined as follows:
    $T=\begin{pmatrix}0&\frac{2ac}{|c|^{2}-1}&0&1\\ 0&1&0&0\\ 1&0&-\frac{2ac}{|c|^{2}-1}&0\\ 0&0&1&0\end{pmatrix}.$
    Since the determinant of $T$ is $1$, we conclude that $T$ is an invertible matrix. The inverse is
    $T^{-1}=\begin{pmatrix}0&0&1&\frac{2ac}{|c|^{2}-1}\\ 0&1&0&0\\ 0&0&0&1\\ 1&-\frac{2ac}{|c|^{2}-1}&0&0\end{pmatrix}.$
    Under this transformation, we verify that $T^{-1}RT=\text{diag}\left(\frac{1}{\bar{c}},\frac{1}{\bar{c}},c,c\right)$. For $n=1$, we obtain:
    $T^{-1}\psi(1)=z\begin{pmatrix}0\\ 0\\ 0\\ \bar{c}\end{pmatrix}+w\begin{pmatrix}\frac{2a}{|c|^{2}-1}\\ 0\\ \bar{c}\\ 0\end{pmatrix},$
    and we compute the $l^{2}$-norm of $T^{-1}\psi$ as:
    $\|T^{-1}\psi\|^{2}=\sum_{n\in\mathbb{N}}\left(\left|\frac{1}{\bar{c}}\right|^{2n}\big{(}|x(n)|^{2}+|y(n)|^{2}\big{)}+|c|^{2n}\big{(}|z(n)|^{2}+|w(n)|^{2}\big{)}\right).$
    Given that the product of the absolute values $|\lambda_{1}||\lambda_{2}|$ is 1 and $|c|<1$, $\|T^{-1}\psi\|^{2}<\infty$ holds if and only if either (i) $z(1)=w(1)=0$, or (ii) $x(1)=y(1)=0$ and $\left|\frac{1}{\bar{c}}\right|<1$.
    1. (a)
      We suppose that $z(1)=w(1)=0$ holds. Since $z(1)=z\bar{c}=0$ and $w(1)=w\bar{c}=0$, and given that $c\neq 0$, we deduce that $z=0$ and $w=0$. Therefore, we determine that the kernel of $H^{\#}-E$ is trivial, i.e., $\ker(H^{\#}-E)=\{0\}$.
    2. (b)
      We suppose that $|c|<1$ and $x(1)=y(1)=0$ hold. In this case, we recognize that $|c|<1$ contradicts the assumption $|c|>1$.
Lemma 6.6.
Suppose that the conditions $c\neq 0$, $b=0$, and $a\neq E$ hold. Under these assumptions, the discriminant $\Delta$ is greater than zero, and the equation $(H^{\#}-E)\psi=0$ has nontrivial $l^{2}$-solutions if and only if $10<|c|<$ 1 and $E=-a\neq 0$. Furthermore, when $10<|c|<$ 1 and $E=-a$, the kernel satisfies:
$\ker(H^{\#}-E)=\langle e_{4}\rangle,$ $\dim(\ker(H^{\#}-E))=1.$
Proof.
We proceed by analyzing the conditions under which nontrivial $l^{2}$-solutions exist and verifying the properties of the kernel.
  1. (1)
    We suppose that $\Delta>0$ holds and that the equation $(H^{\#}-E)\psi=0$ has a nontrivial $l^{2}$-solution. To explore this, we introduce a matrix $T$ defined as follows:
    $T=\begin{pmatrix}0&a_{2}-|c|^{2}&0&a_{1}-|c|^{2}\\ 0&(a-E)c&0&(a-E)c\\ -a_{1}+|c|^{2}&0&-a_{2}+|c|^{2}&0\\ (a-E)\bar{c}&0&(a-E)\bar{c}&0\end{pmatrix}.$
    The determinant of $T$ is $\det T=|c|^{2}(a-E)^{2}\Delta$. Since $\Delta>0$, and given $c\neq 0$ and $a\neq E$, we conclude that $T$ is an invertible matrix. The inverse of $T$ is
    $T^{-1}=\begin{pmatrix}0&0&-\frac{1}{a_{1}-a_{2}}&\frac{|c|^{2}-a_{2}}{\bar{c}(a-E)(a_{1}-a_{2})}\\ -\frac{1}{a_{1}-a_{2}}&-\frac{|c|^{2}-a_{1}}{c(a-E)(a_{1}-a_{2})}&0&0\\ 0&0&\frac{1}{a_{1}-a_{2}}&-\frac{|c|^{2}-a_{1}}{\bar{c}(a-E)(a_{1}-a_{2})}\\ \frac{1}{a_{1}-a_{2}}&\frac{|c|^{2}-a_{2}}{c(a-E)(a_{1}-a_{2})}&0&0\end{pmatrix}.$
    Under this transformation, we verify that $T^{-1}RT=\text{diag}(\lambda_{1},\lambda_{1},\lambda_{2},\lambda_{2})$. We express the transformed solution $T^{-1}\psi(n)$ as the transpose of $(x(n),y(n),z(n),w(n))$. For $n=1$, we obtain:
    $T^{-1}\psi(1)=z\begin{pmatrix}0\\ -\frac{a_{1}}{c(a_{1}-a_{2})}\\ 0\\ \frac{a_{2}}{c(a_{1}-a_{2})}\end{pmatrix}+w\begin{pmatrix}\frac{a^{2}+|c|^{2}-E^{2}-a_{2}}{(a_{1}-a_{2})(a-E)}\\ 0\\ -\frac{a^{2}+|c|^{2}-E^{2}-a_{1}}{(a_{1}-a_{2})(a-E)}\\ 0\end{pmatrix},$
    and we compute the $l^{2}$-norm of $T^{-1}\psi$ as:
    $\|T^{-1}\psi\|^{2}=\sum_{n\in\mathbb{N}}\left(|\lambda_{1}|^{2n}\big{(}|x(1)|^{2}+|y(1)|^{2}\big{)}+|\lambda_{2}|^{2n}\big{(}|z(1)|^{2}+|w(1)|^{2}\big{)}\right).$
    We note that the product of the absolute values $|\lambda_{1}||\lambda_{2}|$ equals to $1$, and we conclude that $\|T^{-1}\psi\|^{2}<\infty$ holds if and only if either (i) $|\lambda_{1}|<1$ and $z(1)=w(1)=0$, or (ii) $|\lambda_{2}|<1$ and $x(1)=y(1)=0$.
    1. (a)
      We suppose that $|\lambda_{1}|<1$ and $z(1)=w(1)=0$ hold. First, we assume that $w=0$. Since $|\lambda_{1}||\lambda_{2}|=1$, we deduce that $a_{2}\neq 0$. From $w(1)=0$, we find that $z=0$, leading to only the trivial solution $\psi=0$. Next, we assume that $w\neq 0$. Since $a_{2}\neq 0$ and $w(1)=w\frac{a^{2}+|c|^{2}-E^{2}-a_{1}}{(a_{1}-a_{2})(a-E)}=0$, we get $a^{2}+|c|^{2}-E^{2}-a_{1}=0$. Additionally, from the relationship between the roots and coefficients of the characteristic polynomial, we have $a_{1}+a_{2}=a^{2}+|c|^{2}+1-E^{2}$, so $a_{1}=1$. Furthermore, since $\lambda_{1}\lambda_{2}=\frac{c}{\bar{c}}$ and $\frac{|c|^{2}}{a^{2}+|c|^{2}-E^{2}}=1$, we obtain $E=-a\neq 0$. Thus, we find $a_{1}=|c|^{2}$, $a_{2}=1$, and $10<|c|=|\lambda_{1}|<$ 1.
    2. (b)
      We suppose that $|\lambda_{2}|<1$ and $x(1)=y(1)=0$ hold. First, we assume that $w=0$. Since $|\lambda_{1}||\lambda_{2}|=1$, we deduce that $a_{1}\neq 0$. From $y(1)=0$, we find that $z=0$, leading to only the trivial solution $\psi=0$. Next, we assume that $w\neq 0$. Since $a_{1}\neq 0$ and $x(1)=w\frac{a^{2}+|c|^{2}-E^{2}-a_{2}}{(a_{1}-a_{2})(a-E)}=0$, we get $a^{2}+|c|^{2}-E^{2}-a_{2}=0$. Additionally, from $a_{1}+a_{2}=a^{2}+|c|^{2}+1-E^{2}$, we obtain $a_{2}=1$. Since $\frac{|c|^{2}}{a^{2}+|c|^{2}-E^{2}}=1$, we again find $E=-a\neq 0$. Thus, we have $a_{1}=1$, $a_{2}=|c|^{2}$, and $10<|c|=|\lambda_{2}|<$ 1.
  2. (2)
    Conversely, we assume that the conditions $10<|c|<$ 1 and $E=-a\neq 0$ hold. Under these conditions, we have $\Delta=(|c|^{2}-1)^{2}$, and since $10<|c|<$ 1, we find $\Delta=-(|c|^{2}-1)>0$. Therefore, we confirm that $a_{1}=|c|^{2}$ and $a_{2}=1$, consistent with the previous case.
Finally, we observe that when $10<|c|<$ 1 and $E=-a$, the kernel is $\ker(H^{\#}-E)=\langle e_{4}\rangle$, with dimension $1$, completing the proof. ∎
Lemma 6.7.
Suppose that the conditions $c\neq 0$ and $b\neq 0$ hold. Under these assumptions, the following holds true:
  1. (1)
    The discriminant $\Delta$ is greater than zero and the equation $(H^{\#}-E)\psi=0$ has nontrivial $l^{2}$-solutions if and only if $10<|c|<$ 1 and $E=\pm\sqrt{a^{2}+|b|^{2}}$.
  2. (2)
    If the conditions $10<|c|<$ 1 and $E=\pm\sqrt{a^{2}+|b|^{2}}$ are satisfied, then the dimension of the kernel is $\dim(\ker(H^{\#}-E))=1$, and one of the eigenvectors is $\frac{a+E}{b}e_{1}+e_{4}$.
  3. (3)
    If the condition $10<|c|<$ 1 holds, then the direct sum of the kernels satisfies $\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}}))=\langle e_{1},e_{4}\rangle$.
Proof.
We proceed by proving each part of the lemma.
  1. (1)
    We suppose that $\Delta>0$ holds and that the equation $(H^{\#}-E)\psi=0$ has a nontrivial $l^{2}$-solution. To analyze this, we introduce a matrix $T$ defined as follows:
    $T=\begin{pmatrix}(a+E)\bar{c}&-a_{2}+|c|^{2}&(a+E)\bar{c}&-a_{1}+|c|^{2}\\ -a_{1}+|c|^{2}&(E-a)c&-a_{2}+|c|^{2}&(E-a)c\\ 0&bc&0&bc\\ b\bar{c}&0&b\bar{c}&0\end{pmatrix}.$
    The determinant of $T$ is $\det T=b^{2}|c|^{2}\Delta$. Since $\Delta>0$, and given $b\neq 0$ and $c\neq 0$, we conclude that $T$ is an invertible matrix. Its inverse is
    $T^{-1}=\begin{pmatrix}0&-\frac{1}{a_{1}-a_{2}}&\frac{E-a}{b(a_{1}-a_{2})}&-\frac{a_{2}-|c|^{2}}{b\bar{c}(a_{1}-a_{2})}\\ \frac{1}{a_{1}-a_{2}}&0&\frac{a_{1}-|c|^{2}}{bc(a_{1}-a_{2})}&-\frac{a+E}{b(a_{1}-a_{2})}\\ 0&\frac{1}{a_{1}-a_{2}}&-\frac{E-a}{b(a_{1}-a_{2})}&\frac{a_{1}-|c|^{2}}{b\bar{c}(a_{1}-a_{2})}\\ -\frac{1}{a_{1}-a_{2}}&0&-\frac{a_{2}-|c|^{2}}{bc(a_{1}-a_{2})}&\frac{a+E}{b(a_{1}-a_{2})}\end{pmatrix}.$
    Under this transformation, we verify that $T^{-1}RT=\text{diag}(\lambda_{1},\lambda_{1},\lambda_{2},\lambda_{2})$. We express the transformed solution $T^{-1}\psi(n)$ as the transpose of
    $(x(n),y(n),z(n),w(n)).$
    For $n=1$, we obtain:
    $T^{-1}\psi(1)=z\begin{pmatrix}0\\ \frac{a_{1}}{c(a_{1}-a_{2})}\\ 0\\ -\frac{a_{2}}{c(a_{1}-a_{2})}\end{pmatrix}+w\begin{pmatrix}\frac{a^{2}+|b|^{2}+|c|^{2}-E^{2}-a_{2}}{b(a_{1}-a_{2})}\\ -\frac{(a+E)a_{1}}{bc(a_{1}-a_{2})}\\ -\frac{a^{2}+|b|^{2}+|c|^{2}-E^{2}-a_{1}}{b(a_{1}-a_{2})}\\ \frac{(a+E)a_{2}}{bc(a_{1}-a_{2})}\end{pmatrix},$
    and we compute the $l^{2}$-norm of $T^{-1}\psi$ as:
    $\|T^{-1}\psi\|^{2}=\sum_{n\in\mathbb{N}}\left(|\lambda_{1}|^{2n}\big{(}|x(1)|^{2}+|y(1)|^{2}\big{)}+|\lambda_{2}|^{2n}\big{(}|z(1)|^{2}+|w(1)|^{2}\big{)}\right).$
    We note that the product of the absolute values $|\lambda_{1}||\lambda_{2}|$ equals 1, and we conclude that $\|T^{-1}\psi\|^{2}<\infty$ holds if and only if either (i) $|\lambda_{1}|<1$ and $z(1)=w(1)=0$, or (ii) $|\lambda_{2}|<1$ and $x(1)=y(1)=0$.
    1. (a)
      We suppose that $|\lambda_{1}|<1$ and $z(1)=w(1)=0$ hold. From $z(1)=0$, we deduce that $a_{1}=a^{2}+|b|^{2}+|c|^{2}-E^{2}$. By using the relationship between the roots and coefficients of the characteristic polynomial, we have $a_{1}+a_{2}=a^{2}+|b|^{2}+|c|^{2}+1-E^{2}$, so $a_{2}=1$ and $a_{1}=|c|^{2}$. Additionally, since $a_{1}a_{2}=\frac{|c|^{2}}{a^{2}+|b|^{2}+|c|^{2}-E^{2}}$, we find $E=\pm\sqrt{a^{2}+|b|^{2}}$. We also note that $|c|=|\lambda_{1}|<1$, and from $w(1)=-\frac{zb-w(a+E)}{bc(a_{1}-a_{2})}=0$, we obtain $z=\frac{a+E}{b}w$.
    2. (b)
      We suppose that $|\lambda_{2}|<1$ and $x(1)=y(1)=0$ hold. From $x(1)=0$, we deduce that $a_{2}=a^{2}+|b|^{2}+|c|^{2}-E^{2}$. By using $a_{1}+a_{2}=a^{2}+|b|^{2}+|c|^{2}+1-E^{2}$, we find $a_{1}=1$ and $a_{2}=|c|^{2}$. Since $a_{1}a_{2}=\frac{|c|^{2}}{a^{2}+|b|^{2}+|c|^{2}-E^{2}}$, it follows that $E=\pm\sqrt{a^{2}+|b|^{2}}$. We note that $|c|=|\lambda_{2}|<1$, and from $w(1)=-\frac{zb-w(a+E)}{bc(a_{1}-a_{2})}=0$, we obtain $z=\frac{a+E}{b}w$.
  2. (2)
    Conversely, we assume that the conditions $10<|c|<$ 1 and $E=\pm\sqrt{a^{2}+|b|^{2}}$ hold. Under these conditions, we have $\Delta=(|c|^{2}-1)^{2}$, and since $10<|c|<$ 1, we find that $\Delta=-(|c|^{2}-1)>0$. Thus, we confirm that $a_{1}=|c|^{2}$ and $a_{2}=1$, ensuring that $\Delta>0$ and that $(H^{\#}-E)\psi=0$ has nontrivial $l^{2}$-solutions.
Next, we verify the remaining claims. Since we established that $z=\frac{a+E}{b}w$, we compute $\psi(1)$ as:
$\psi(1)=w\begin{pmatrix}\frac{a+E}{b}\bar{c}\\ 0\\ 0\\ \bar{c}\end{pmatrix}=w\bar{c}\left(\frac{a+E}{b}e_{1}(1)+e_{4}(1)\right).$
Given $a_{1}=|c|^{2}$ and $a_{1}-a_{2}=|c|^{2}-1=-\Delta$, we derive the general form of $\psi(n)$ as follows:
$\psi(n)$ $=R^{n-1}\psi(1)=T(T^{-1}RT)^{n-1}T^{-1}\psi(1)$
$=T\text{diag}\left(c^{n-1},c^{n-1},\left(\frac{1}{\bar{c}}\right)^{n-1},\left(\frac{1}{\bar{c}}\right)^{n-1}\right)T^{-1}\psi(1).$
Applying $T^{-1}\psi(1)$, we simplify:
$T^{-1}\psi(1)=w\begin{pmatrix}\frac{|c|^{2}-1}{b(a_{1}-a_{2})}\\ 0\\ 0\\ 0\end{pmatrix},$
and thus:
$\psi(n)$ $=wT\begin{pmatrix}c^{n-1}\frac{|c|^{2}-1}{b(a_{1}-a_{2})}\\ 0\\ 0\\ 0\end{pmatrix}$
$=w\bar{c}\frac{|c|^{2}-1}{a_{1}-a_{2}}\left(\frac{a+E}{b}e_{1}(n)+e_{4}(n)\right)=w\bar{c}\left(\frac{a+E}{b}e_{1}(n)+e_{4}(n)\right).$
Therefore, we identify one eigenvector with $\frac{a+E}{b}e_{1}+e_{4}$, and since the kernel is one-dimensional (due to the single free parameter $w$), we have $\dim(\ker(H^{\#}-E))=1$. Finally, we confirm that the direct sum $\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}}))=\langle e_{1},e_{4}\rangle$, as each kernel is spanned by a linearly independent eigenvector, completing the proof. ∎
Proposition 6.8.
The following holds true.
  1. (1)
    The equation $(H^{\#}-E)\psi=0$ has a nontrivial $l^{2}$-solution if and only if one of the following conditions holds:
    1. (a)
      The parameters satisfy $c=0$ and $E=\pm\sqrt{a^{2}+|b|^{2}+1}$. Under this condition, the dimension of the kernel is infinite, i.e., $\dim(\ker(H^{\#}-E))=\infty$.
    2. (b)
      The parameters satisfy $b=0$, $E=a=0$, and $10\leq|c|<$ 1. Under this condition, the dimension of the kernel is 2, i.e., $\dim(\ker(H^{\#}-E))=2$.
    3. (c)
      The parameters satisfy $10\leq|c|<$ 1 and $E=\sqrt{a^{2}+|b|^{2}}\neq 0$. Under this condition, the dimension of the kernel is 1, i.e., $\dim(\ker(H^{\#}-E))=1$.
  2. (2)
    The basis vectors $e_{1}$ and $e_{4}$ induce the following linear isomorphisms:
    1. (a)
      We suppose that the conditions $b=0$ and $E=a=0$ hold. Under these conditions, we have $\ker(H^{\#})\cong\mathbb{C}^{2}$.
    2. (b)
      We suppose that the conditions $10\leq|c|<$ 1 and $E=\sqrt{a^{2}+|b|^{2}}\neq 0$ hold. Under these conditions, we have $\ker(H^{\#}-\sqrt{a^{2}+|b|^{2}})\oplus\ker(H^{\#}-(-\sqrt{a^{2}+|b|^{2}}))\cong\mathbb{C}^{2}$.
  3. (3)
    Under the linear isomorphisms described above, we can express the representation matrix of $H^{\#}$ as follows:
    $\begin{pmatrix}a&\bar{b}\\ b&-a\end{pmatrix}=\text{Re}(b)\sigma_{1}+\text{Im}(b)\sigma_{2}+a\sigma_{3},$
    where $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ are Pauli matrices
    $\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\sigma_{2}=\begin{pmatrix}0&-\sqrt{-1}\\ \sqrt{-1}&0\end{pmatrix},\quad\sigma_{3}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$
Proof.
We derive this proposition from the results of the previous lemmas (6.1, 6.2, 6.3, 6.4, 6.5, 6.6, 6.7) and straightforward calculations. ∎